### v1

# Completeness of closed Kleinian flat Pseudo-Riemannian Manifolds of Signature (2,2)

Abstract.
Let $\mathbb{R}^{2,2}$ denote the model space of flat pseudo-Riemannian manifolds of signature $(2,2)$. We prove that the only domain divisible by a discrete subgroup of the isometry group of $\mathbb{R}^{2,2}$ is $\mathbb{R}^{2,2}$ itself. In the Kleinian setting, this provides the first completeness theorem of closed flat pseudo-Riemannian manifolds beyond the Euclidean and Lorentzian cases.
Along the proof, we show two results of independent interest. The first is a geometric reduction for certain divisible domains of affine space. The second concerns the existence of syndetic hulls in semidirect products $R\ltimes G$, where $G$ is a homothety Lie group. This construction generalizes earlier constructions in affine geometry due to Carrière and Dal’bo.
1. Introduction
Since Felix Klein’s Erlangen program, geometric structures have played an important role in geometry and topology. From this viewpoint, a geometry is determined by a model space $X$ together with a transitive and effective action of a Lie group $G$; a manifold $M$ carries a $(G,X)$-structure if it admits an atlas of charts with values in $X$ whose transition maps are restrictions of elements of $G$. Affine structures have gained particular attention in this context. An affine manifold is a manifold endowed with an $(\operatorname{Aff}(\mathbb{R}^{d}),\mathbb{R}^{d})$-structure, where $\operatorname{Aff}(\mathbb{R}^{d})={\operatorname{GL}}_{d}(\mathbb{R})\ltimes\mathbb{R}^{d}$ is the group of affine transformations of $\mathbb{R}^{d}$. Many fundamental questions on affine structures remain open, although they have been extensively studied; see for instance [Fri80, FGH81, Fri82, FG83, GH84, Fri86, GH86, Car89, GLM09, DGK16, Kli17, AMS20]. One of the most important questions, raised by Markus in 1960, concerns the completeness of special affine manifolds. An affine manifold is said to be complete if it is the quotient of $\mathbb{R}^{d}$ by a discrete group of affine transformations acting freely and properly discontinuously. The Markus conjecture asserts that a closed affine manifold with parallel volume, that is, whose linear holonomy lies in ${\operatorname{SL}}_{d}(\mathbb{R})$, is complete. In the Markus conjecture, the assumption of having parallel volume is necessary. Indeed, the so-called Hopf manifolds are known examples of incomplete closed affine manifolds. A Hopf manifold is the quotient of $\mathbb{R}^{d}\setminus\{0\}$ by the discrete group generated by $\lambda{\operatorname{Id}}$ for some $\lambda>1$. Topologically, $d$-dimensional Hopf manifolds are diffeomorphic to $\mathbb{S}^{1}\times\mathbb{S}^{d-1}$. This class of manifolds belongs to the family of Kleinian affine manifolds. These are manifolds that are finite covers of $\Omega/\Gamma$, where $\Omega\subset\mathbb{R}^{d}$ is a domain, i.e. a nonempty connected open set, and $\Gamma\leq\operatorname{Aff}(\mathbb{R}^{d})$ is a discrete group acting freely, properly discontinuously, and cocompactly on $\Omega$ (see Definition 2.1). When the linear part of $\Gamma$ lies in ${\operatorname{SL}}_{d}(\mathbb{R})$, one may formulate a Kleinian version of Markus’ conjecture:
Kleinian Markus conjecture. Let $\Omega$ be a domain of $\mathbb{R}^{d}$ and let $\Gamma\leq{\operatorname{SL}}_{d}(\mathbb{R})\ltimes\mathbb{R}^{d}$ be a discrete group acting freely, properly discontinuously, and cocompactly on $\Omega$. Then $\Omega=\mathbb{R}^{d}$.
The conjecture is solved in dimension two: it follows from the completeness of special affine surfaces, see for instance, [Ben60]. In higher dimensions $(d\geq 3)$, the question remains widely open. There are partial results under additional hypotheses: one may assume that $\Gamma$ preserves a geometric structure stronger than a parallel volume form. For example, if $\Gamma$ preserves a pseudo-Riemannian metric of signature $(p,q)$, equivalently, if $\Gamma$ is a discrete subgroup of ${\operatorname{SO}}(p,q)\ltimes\mathbb{R}^{p+q}$, then the Kleinian Markus conjecture holds in the Riemannian case $q=0$ by the classical Hopf–Rinow theorem, and in the Lorentzian case $q=1$ by Carrière’s completeness theorem for closed flat Lorentzian manifolds [Car89]. Outside these signatures, the problem is largely open; the main contribution of this paper is a positive answer in signature $(2,2)$, which provides the first evidence for the validity of completeness of closed flat pseudo-Riemannian manifolds manifolds of non-Lorentzian signature.
Another geometric condition is provided when $\Gamma$ preserves a pseudo-Hermitian structure, i.e., when $\Gamma$ is a discrete subgroup of $\mathrm{U}(p,q)\ltimes\mathbb{C}^{\,p+q}$. In this setting, the conjecture is known when $q=0$ (the Riemannian case) and when $q=1$ by a result of Tholozan [Tho15]. The latter structures are known in the literature as Hermite-Lorentz manifolds (see for instance [AZ16, Bar20]). In the same paper, Tholozan also obtained a positive answer when $\Gamma$ is a discrete subgroup of ${\operatorname{SO}}(3,\mathbb{C})\ltimes\mathbb{C}^{3}$ acting on $\mathbb{C}^{3}$, which is the model of flat holomorphic Riemannian manifolds. In a different direction, Jo–Kim [JK21] proved the Kleinian Markus conjecture for convex domains $\Omega\subset\mathbb{R}^{d}$ when $d\leq 5$.
There are also results of a different flavor that that address the completeness problem under algebraic assumptions on $\Gamma$, for instance when $\Gamma$ is abelian or nilpotent; see [Smi77, FGH81, GH86, Fri86]. We return to some of these results in more detail in the preliminaries (Section 2).
1.1. Main result
Let $\mathbb{R}^{2,2}$ denote the affine space $\mathbb{R}^{4}$ equipped with a nondegenerate quadratic form of signature $(2,2)$. The group of orientation-preserving isometries of $\mathbb{R}^{2,2}$ is ${\operatorname{SO}}(2,2)\ltimes\mathbb{R}^{2,2}$. In particular, $\mathbb{R}^{2,2}$ is a homogeneous flat pseudo-Riemannian space of signature $(2,2)$. Our first main result is the following.
Theorem A.
Let $\Omega$ be a domain of $\mathbb{R}^{2,2}$ and let $\Gamma\leq{\operatorname{SO}}(2,2)\ltimes\mathbb{R}^{2,2}$ be a discrete group acting freely, properly discontinuously, and cocompactly on $\Omega$. Then $\Omega=\mathbb{R}^{2,2}$.
In other words, Kleinian $({\operatorname{SO}}(2,2)\ltimes\mathbb{R}^{2,2},\mathbb{R}^{2,2})$-closed manifolds are complete. This provides the first result toward the Kleinian Markus conjecture in dimension $4$, which remains open. Note that our result does not assume any additional topological conditions on the domain $\Omega$.
Carrière [Car89] proved Markus’ conjecture when the linear part of $\Gamma$ has discompacity $1$ (for instance, when it is contained in ${\operatorname{SO}}(n,1)$). Roughly speaking, discompacity measures the number of contracting directions of an ellipsoid under the action of the linear part of $\Gamma$. The discompacity of a reductive group is in general bigger than the (real) rank of the group; for instance, the irreducible embedding of ${\operatorname{SL}}_{2}(\mathbb{R})$ into ${\operatorname{SL}}_{d}(\mathbb{R})$ has discompacity $\lfloor d/2\rfloor$. Discompacity $1$ can therefore be seen as a strong rank-one assumption. Tholozan later extended Carrière’s argument to prove completeness for Kleinian $(\mathrm{U}(n,1)\ltimes\mathbb{C}^{\,n+1},\mathbb{C}^{\,n+1})$-affine manifolds; his approach relies on the fact that $\mathrm{U}(n,1)$ has complex discompacity $1$, thereby going beyond the strictly real framework of [Car89]. However, both these results are restricted to manifolds whose linear holonomy is contained in a rank-one simple Lie group (and even satisfying the stronger condition of discompacity $1$). Theorem A is the first result related to Kleinian Markus conjecture that addresses manifolds with no further constraint than a linear holonomy in a higher-rank simple Lie group. In particular, it does not follow either from Carrière’s result (the group ${\operatorname{SO}}(2,2)$ has discompacity $2$) or from Tholozan’s result (the standard embedding of $\mathrm{U}(n,1)$ into ${\operatorname{SO}}(2n,2)$ preserves a complex structure on $\mathbb{C}^{\,n+1}\cong\mathbb{R}^{2n,2}$, which imposes additional rigidity).
1.2. Strategy of the proof
We now provide an overview of the techniques used in the proof of Theorem A. We may always replace $\Gamma$ by a finite-index subgroup, this allow us to assume that the linear part of $\Gamma$ is contained in the identity component ${\operatorname{SO}}_{0}(2,2)$ of ${\operatorname{SO}}(2,2)$. More generally, we will reason in terms of virtual algebraic properties of $\Gamma$, that is, properties considered up to finite index.
1.2.1. Reduction result
The first main idea in the proof of Theorem A is the following. Let $\Omega\subset\mathbb{R}^{2,2}$ be as in Theorem A. If $\Omega\neq\mathbb{R}^{2,2}$, then it is foliated by parallel isotropic planes in $\mathbb{R}^{2,2}$. This fact is established through the more general Proposition B below.
Let $G\leq{\operatorname{SL}}(d,\mathbb{R})$ be a semisimple Lie subgroup, and fix a Cartan decomposition $G=KA^{+}K$, where $K\leq G$ is a maximal compact subgroup and $A^{+}$ is the exponential of a closed positive Weyl chamber (see Section 3.1). For a subgroup $G^{\prime}\leq G$, denote by $A^{+}(G^{\prime})$ the set of Cartan projections of elements of $G^{\prime}$ with respect to this decomposition. Fix once and for all a Euclidean norm $|\cdot|$ on $\mathbb{R}^{d}$ that is invariant under the maximal compact subgroup $K$. We say that $A^{+}(G^{\prime})$ does not contract a subspace $F\subset\mathbb{R}^{d}$ with respect to $|\cdot|$ if $|g\cdot x|\geq|x|$ for all $g\in A^{+}(G^{\prime})$ and all $x\in F$.
Fix $p\in\llbracket 1,d\rrbracket$, and consider $\operatorname{Gr}_{p}(\mathbb{R}^{d})$, the Grassmannian of $p$-dimensional subspaces. We say that $E,F\in\operatorname{Gr}_{p}(\mathbb{R}^{d})$ are transverse if $E+F=\mathbb{R}^{d}$ (in particular $p\geq\tfrac{d}{2}$). We prove the following result.
Proposition B.
Let $\Omega$ be a domain in $\mathbb{R}^{d}$, different from $\mathbb{R}^{d}$, and let $\Gamma\leq{\operatorname{SL}}_{d}(\mathbb{R})\ltimes\mathbb{R}^{d}$ be a discrete group acting freely, properly discontinuously, and cocompactly on $\Omega$. Suppose that the linear part $L(\Gamma)$ of $\Gamma$ is contained in a semisimple Lie group $G\leq{\operatorname{SL}}_{d}(\mathbb{R})$ such that:
  1. (1)
    there exists $\frac{d}{2}\leq p\leq d$ such that $G$ preserves an orbit $\mathcal{F}\subset\operatorname{Gr}_{p}(\mathbb{R}^{d})$ whose closure consists of pairwise transverse subspaces;
  2. (2)
    there exists $H\in\mathcal{F}$ such that $A^{+}(L(\Gamma))$ does not contract $H$.
Then there exists $H^{\prime}$ in the closure of $\mathcal{F}$ that is $L(\Gamma)$-invariant, and $\Omega$ is foliated by affine subspaces parallel to $H^{\prime}$.
This proposition generalizes a result of Tholozan [Tho15, Proposition 2.4], where the statement is proved in the case $G=\mathrm{U}(n,1)$ and $\mathcal{F}$ is the orbit of a complex hyperplane. In our case, the set of isotropic planes in $\mathbb{R}^{2,2}$ has two connected components, each of which is an orbit under the identity component ${\operatorname{SO}}_{0}(2,2)$; moreover, the elements of each orbit are pairwise transverse (see Proposition 3.7). Let us briefly explain this well known fact. Write $\mathbb{R}^{2,2}=\mathbb{R}^{2}\oplus\mathbb{R}^{2}$, endowed with the nondegenerate quadratic form of signature $(2,2)$ given by $g\oplus(-g)$, where $g$ is the standard Euclidean metric. The graph of any linear isometry $A\in O(2)$ is a totally isotropic plane in $\mathbb{R}^{2,2}$. One can show that the set of isotropic planes in $\mathbb{R}^{2,2}$ is exactly $O(2)$; in particular, it has two connected components. The key observation is that if $A,B\in{\operatorname{SO}}(2)$ are distinct elements, then their graphs intersect only at $0\in\mathbb{R}^{2,2}$ and so they are transverse. Indeed, their intersection coincides with the set of fixed vectors of $AB^{-1}$, and a nontrivial element of ${\operatorname{SO}}(2)$ has no nonzero fixed vectors. This property no longer holds in ${\operatorname{SO}}(n)$ for $n>2$, and it is precisely at this point that the signature $(2,2)$ plays a crucial role. Nevertheless, we believe that Proposition B is of independent interest and may apply in other settings, which deserve further investigation. For instance, one may consider the case where $G={\operatorname{SO}}(4,\mathbb{C})$, as well as the case where $G$ is a rank-one irreducible simple Lie subgroup of ${\operatorname{SL}}_{d}(\mathbb{R})$. These cases are part of an ongoing project of the three authors.
Having established this, Proposition B applies in our setting and yields a foliation of any domain $\Omega\subset\mathbb{R}^{2,2}$ as in Theorem A by parallel isotropic planes. The linear part of $\Gamma$ stabilizes an isotropic plane in $\mathbb{R}^{2,2}$, so that, up to conjugacy, $\Gamma$ is contained in the group ${\operatorname{GL}}_{2}^{+}(\mathbb{R})\ltimes N$, where ${\operatorname{GL}}_{2}^{+}(\mathbb{R})$ denotes the identity component of ${\operatorname{GL}}_{2}(\mathbb{R})$ and $N$ is a step-two nilpotent group (see Lemma 3.11). This produces two natural projections:
  1. (1)
    $p:{\operatorname{GL}}_{2}^{+}(\mathbb{R})\ltimes N\to{\operatorname{GL}}_{2}^{+}(\mathbb{R})$, the natural projection onto the ${\operatorname{GL}}_{2}^{+}(\mathbb{R})$ factor;
  2. (2)
    $q:{\operatorname{GL}}_{2}^{+}(\mathbb{R})\ltimes N\to{\operatorname{GL}}_{2}^{+}(\mathbb{R})\ltimes\mathbb{R}^{2}$, the projection modulo the isotropic foliation. If $P_{0}$ denotes the linear part of the isotropic foliation, then $\Omega$ fibers over a domain $\widehat{\Omega}\subset\mathbb{R}^{2}\cong\mathbb{R}^{2,2}/P_{0}$, which is preserved by $q(\Gamma)$.
In the language of foliated geometric structures, the quotient $M=\Omega/\Gamma$ carries a transversal $({\operatorname{GL}}_{2}(\mathbb{R})\ltimes\mathbb{R}^{2},\mathbb{R}^{2})$-foliation; for background on foliated geometric structures, see [Blu79, Eps83, Car84, Thu22].
1.2.2. Quotient geometry
Despite the last reduction, the group ${\operatorname{GL}}_{2}^{+}(\mathbb{R})\ltimes N$ still has discompacity $2$, and therefore Carrière’s Theorem [Car89] cannot be applied directly. A natural strategy is then to study the projection $p(\Gamma)$ onto the ${\operatorname{GL}}_{2}^{+}(\mathbb{R})$ factor. To this end, we observe that the $N$ factor belongs to the class of homothety Lie groups (see Definition 2.8 and Lemma 3.12). This class includes, for instance, abelian Lie groups, Heisenberg groups, or more generally Carnot groups, which are higher-step nilpotent generalizations of Heisenberg groups. This observation leads us to establish the second main result of the paper, which goes beyond the context of Theorem A.
Theorem C.
Let $G$ be a homothety Lie group, and let $G_{\theta}=R\ltimes_{\theta}G$ be a semidirect product, where $R$ is a linear Lie group and $\theta$ commutes with a nontrivial homothety of $G$. Let $\pi\colon G_{\theta}\to R$ be the natural projection. Let $\Gamma\leq G_{\theta}$ be a discrete subgroup, and let $H=\overline{\pi(\Gamma)}$. Set $\Gamma_{nd}:=\Gamma\cap\pi^{-1}(H^{\circ})$. Then $\Gamma_{nd}$ admits a nilpotent syndetic hull $S$ in $G_{\theta}$ with the property that $\overline{\pi(\Gamma_{nd})}=\overline{\pi(N)}=H^{\circ}$.
A syndetic hull of $\Gamma_{nd}$ is a closed connected subgroup containing $\Gamma_{nd}$ as a uniform lattice. Theorem C is applied at several points in the proof of Theorem A. In particular, it implies that $\overline{p(\Gamma)}^{\circ}$ is a nilpotent subgroup of ${\operatorname{GL}}_{2}^{+}(\mathbb{R})$. This observation plays a key role when the group $p(\Gamma)$ is not discrete; see, for instance, Lemma 5.22.
Theorem C generalizes a theorem of Carrière–Dal’bo [CD89], who proved an analogous result for the affine group ${\operatorname{GL}}_{d}(\mathbb{R})\ltimes\mathbb{R}^{d}$. Another special case was proved recently in [HKMZ25, Theorem 1.4]. We believe that Theorem C is of independent interest and may be applied in different contexts.
Having established this, we proceed with the proof of Theorem A according to the cohomological dimension of the discrete abelian group $\Gamma\cap{\operatorname{Ker}}(q)$. Since $\Gamma\cap{\operatorname{Ker}}(q)$ preserves each leaf of the isotropic foliation and acts freely and properly discontinuously on it, its cohomological dimension is at most $2$. We use results that ensure completeness under algebraic hypotheses. The first is due to Fried–Goldman–Hirsch [FGH81], which states that if $\Gamma$ is nilpotent and the affine structure admits a parallel volume form, then the affine structure is complete. Another result of Goldman–Hirsch [GH86] treats the solvable case. The difficult part is to reach a situation where the latter result applies; this requires, in particular, unimodularity arguments and a careful analysis of syndetic hulls provided by Theorem C. These arguments are developed in Sections 5-6.
1.3. Organization of the paper
In Section 2, we collect some preliminary results on closed flat affine manifolds. In Section 3, we prove Proposition B, which provides the reduction result. Section 4 is devoted to the study of domains in the two-dimensional affine space that are invariant under particular one-parameter groups. In Section 5, we prove completeness in the case where $\Gamma\cap{\operatorname{Ker}}(q)$ is nontrivial, while Section 6 treats the case where $\Gamma\cap{\operatorname{Ker}}(q)$ is trivial. Theorem C is proved in Section 7 and is part of the PhD thesis of the third author. The proof of Theorem C and can be read independently of Sections 26. Finally, the appendix contains auxiliary results on the structure of certain abelian subgroups of ${\operatorname{GL}}_{2}^{+}(\mathbb{R})\ltimes N$ that are used at various points in the paper.
1.4. Acknowledgments
We would like to thank Andrea Seppi for helpful comments on an earlier version of this manuscript.
2. Preliminaries on affine geometry
We start by reviewing some basic definitions about affine geometry and collect the results we use throughout the paper. Most of the statements below are standard; we refer the reader to the survey [DDGS20] for a more detailed exposition on the topic.
2.1. Affine manifolds
An affine manifold of dimension $d$ is a manifold $M$ endowed with an atlas of charts with values in $\mathbb{R}^{d}$ whose transition maps are restrictions of elements of $\operatorname{Aff}(\mathbb{R}^{d})={\operatorname{GL}}_{d}(\mathbb{R})\ltimes\mathbb{R}^{d}$, the group of affine transformations of $\mathbb{R}^{d}$. An important result in the theory is that the data of an affine manifold $M$ is equivalent to the data of a developing map $\operatorname{dev}:\widetilde{M}\to\mathbb{R}^{d}$, which is a local diffeomorphism equivariant with respect to the holonomy representation $\operatorname{hol}:\pi_{1}(M)\to\operatorname{Aff}(\mathbb{R}^{d})$. That is, for any $x\in\widetilde{M}$ and $\gamma\in\pi_{1}(M)$,
$\operatorname{dev}(\gamma\cdot x)=\operatorname{hol}(\gamma)\,\operatorname{dev}(x).$
We call the image $\operatorname{hol}(\pi_{1}(M))$ the affine holonomy. The pair $(\operatorname{dev},\operatorname{hol})$ is defined only up to the action of $\operatorname{Aff}(\mathbb{R}^{d})$, where $\operatorname{Aff}(\mathbb{R}^{d})$ acts by conjugation on the holonomy representation and by post-composition on the developing map.
An affine manifold is said to be complete if $\operatorname{dev}$ is a diffeomorphism onto $\mathbb{R}^{d}$. In this case $M$ is a quotient of $\mathbb{R}^{d}$ by a discrete subgroup $\Gamma\leq\operatorname{Aff}(\mathbb{R}^{d})$ acting freely and properly discontinuously on $\mathbb{R}^{d}$. A particularly nice class of affine manifold is given by Kleinian structures. Following [KP06], we make the following definition.
Definition 2.1.
[…]
We say that a domain $\Omega\subset\mathbb{R}^{d}$ is divisible by a discrete subgroup $\Gamma\leq\operatorname{Aff}(\mathbb{R}^{d})$ if $\Gamma$ acts freely, properly discontinuously, and cocompactly on $\Omega$; we then say that $\Omega$ is divided by $\Gamma$. The quotient $\Omega/\Gamma$ is a Kleinian affine manifold. Observe that if $M=\Omega/\Gamma$ is a complete affine manifold, then $\Omega=\mathbb{R}^{d}$. Indeed, in this case the developing map is the covering $\operatorname{dev}:\widetilde{M}\to\Omega\subset\mathbb{R}^{d}$, and the holonomy representation $\operatorname{hol}:\pi_{1}(M)\to\operatorname{Aff}(\mathbb{R}^{d})$ is the one associated to this covering, so that $\operatorname{hol}(\pi_{1}(M))=\Gamma$. Completeness of $M$ implies, in particular, that $\operatorname{dev}$ is surjective, and hence $\Omega=\mathbb{R}^{d}$.
2.2. Algebraic and cohomological constraints
We recall some algebraic obstructions for the affine holonomy of an affine manifold admitting a parallel volume form, i.e., an affine manifold whose holonomy lies in ${\operatorname{SL}}_{d}(\mathbb{R})\ltimes\mathbb{R}^{d}$. The first result states that the affine holonomy is irreducible.
Theorem 2.2.
[GH86, Thm. p. 182 and Cor. 2.5] Let $M$ be a closed affine manifold endowed with a parallel volume form. Then the affine holonomy does not preserve any proper algebraic subset of $\mathbb{R}^{d}$. In particular, it does not preserve any proper affine subspace of $\mathbb{R}^{d}$.
The next results deal with the cohomological dimension of groups. For a torsion-free group $\Gamma$, we denote its integral cohomological dimension by $\operatorname{cd}(\Gamma)$. We record below the basic facts that will be used throughout the paper and refer the reader to [Bro82, VIII.2] for further details.
  1. (1)
    Let $\Gamma$ be a torsion-free group acting properly discontinuously and freely on a contractible manifold $X$. Then $\operatorname{cd}(\Gamma)\leq\dim(X)$, with equality if and only if the action is cocompact; see [Bro82, VIII. Proposition 8.1].
  2. (2)
    If $1\to\Gamma^{\prime}\to\Gamma\to\Gamma^{\prime\prime}\to 1$ is a short exact sequence of torsion-free groups, then
    $\operatorname{cd}(\Gamma)\leq\operatorname{cd}(\Gamma^{\prime})+\operatorname{cd}(\Gamma^{\prime\prime}),$
    see [Bro82, VIII. Proposition 2.4].
Recall that, according to Selberg’s Lemma [Sel60], any finitely generated subgroup of a linear group is virtually torsion-free. In the situations that concern us, we may therefore assume without loss of generality that the discrete linear groups under consideration are torsion-free.
We now recall the following result, which will be used in the sequel.
Theorem 2.3 ([GH84, Cor. 2.11]).
Let $\Omega\subset\mathbb{R}^{d}$ be an open set divided by a discrete subgroup $\Gamma\leq{\operatorname{SL}}(d,\mathbb{R})\ltimes\mathbb{R}^{d}$. Then $\operatorname{cd}(\Gamma)\geq d$.
For instance, this implies that a free group cannot divide an open set in $\mathbb{R}^{d}$ when $d\geq 2$, since the cohomological dimension of such group is $1$. It is worth noting that if $\Omega$ were assumed to be contractible, then the conclusion would be immediate, as one would have $\operatorname{cd}(\Gamma)=d$. However, the theorem makes no such assumption on the open set, which makes the result useful later on.
2.3. Completeness results
As indicated in the introduction, Markus’ conjecture states that a closed affine manifold with parallel volume form should be complete. Although the conjecture remains widely open, there are completeness results under additional assumptions on the affine holonomy. The following result ensures completeness in the case where the affine holonomy is nilpotent.
Theorem 2.4 ([FGH81]).
Let $M$ be a closed affine manifold with parallel volume form. If the affine holonomy is nilpotent, then $M$ is complete.
Another algebraic condition on the affine holonomy implying completeness is the following:
Theorem 2.5.
[GH86, Theorem 3.5] Let $M$ be a closed affine manifold with parallel volume form. Assume that the affine holonomy is solvable with cohomological dimension equal to the dimension of $M$. Then $M$ is complete.
2.4. Syndetic hulls
We finish this preliminaries section by recalling the notion of a syndetic hull, an important tool in the proof of our main result.
Definition 2.6.
Let $G$ be a Lie group and let $\Gamma$ be a discrete subgroup. A syndetic hull of $\Gamma$ in $G$ is a closed connected Lie subgroup $S\leq G$ containing $\Gamma$ as a uniform lattice, that is, the quotient $\S\Gamma\backslash$ S is compact.
Observe that if $S$ is a syndetic hull of a torsion-free group $\Gamma$, then $\dim(S)\geq\operatorname{cd}(\Gamma)$. Indeed, if $K$ denotes the maximal compact subgroup of $S$, then $\Gamma$ acts properly, cocompactly and freely on the contractible space $S/K$, and therefore $\operatorname{cd}(\Gamma)=\dim(S/K)=\dim(S)-\dim(K)$. We recall the following existence theorem for syndetic hulls.
Theorem 2.7.
[FG83, Section 1.6] Let $G$ be a linear algebraic Lie group and let $\Gamma\leq G$ be a virtually solvable discrete subgroup. Then virtually $\Gamma$ admits a solvable syndetic hull $S\leq G$. Moreover, the Zariski closure of $S$ coincides with the Zariski closure of $\Gamma$ in $G$.
In simply connected nilpotent groups $G$, the existence of a syndetic hull is due to Malcev [Mal49, Rag72], and it is unique; this is called the Malcev closure. Beyond the nilpotent and solvable cases, there is another interesting setting that guarantees the existence of a syndetic hull, provided by Theorem C. For this, we introduce the following definition.
Definition 2.8 (Homothety Lie group).
Let $G$ be a real Lie group and $\Psi\in{\sf{Aut}}(G)$ an automorphism. We say that $\Psi$ is a homothety if the differential $d\Psi$ or $d\Psi^{-1}$ at the identity is diagonalizable over $\mathbb{R}$ and all its eigenvalues have absolute value greater than $1$. A Lie group that admits a homothety is called a homothety Lie group.
The reader interested in the proof of Theorem C may refer directly to Section 7.
3. Reduction of the flat affine geometry
The goal of this section is to prove the following proposition:
Proposition 3.1.
Let $\Omega\subset\mathbb{R}^{2,2}$ be a proper domain, divided by a subgroup $\Gamma\leq{\operatorname{SO}}_{0}(2,2)\ltimes\mathbb{R}^{2,2}$. Then the linear part of $\Gamma$ preserves a $2$-dimensional totally isotropic plane $P_{0}\subset\mathbb{R}^{2,2}$, and $\Omega$ is foliated by affine translates of $P_{0}$.
3.1. Cartan projection
An important feature of the proof of Proposition 3.1 is the Cartan projection. Let $G$ be a semisimple Lie group, and denote by $\mathfrak{g}$ its Lie algebra. Let $K\leq G$ be a maximal compact subgroup and $\mathfrak{h}$ be the orthogonal of the Lie algebra $\mathfrak{k}$ of $K$ in $\mathfrak{g}$} for the Killing form. A Cartan subspace $\mathfrak{a}$ of $\mathfrak{g}$} is a maximal abelian subalgebra of $\mathfrak{h}$.
The restricted Weyl group $W$ of $G$ is the quotient of the normalizer of $\mathfrak{a}$ in $K$ (for the adjoint action) by the centralizer of $\mathfrak{a}$ in $K$. A fundamental system of $\mathfrak{a}$ for the action of $W$ is called a closed Weyl chamber. It is a closed convex cone with nonempty interior. Given such a closed Weyl chamber $\overline{\mathfrak{a}}^{+}$, we denote by $A^{+}$ the exponential of $\overline{\mathfrak{a}}^{+}$ ; the group $G$ then admits the following Cartan decomposition
$G=KA^{+}K.$
Given an element $g\in G$, we can thus write $g=kak^{\prime}$ according to this decomposition. The elements $k,k^{\prime}\in K$ are not unique, but the element $a\in A^{+}$ is. The Cartan projection of $g$ is the unique element $X\in\overline{\mathfrak{a}}^{+}$ such that $a=\exp(X)$.
3.2. General reduction result
For a semisimple Lie subgroup $G\leq{\operatorname{SL}}(d,\mathbb{R})$, we write $G=KA^{+}K$ for a chosen Cartan decomposition of $G$ (in the notation of Section 3.1 above). For a subgroup $G^{\prime}\leq G$, we denote by $A^{+}(G^{\prime})$ the set of Cartan projections of elements of $G^{\prime}$ in this decomposition.
Fix once and for all a Euclidean norm $|\cdot|$ on $\mathbb{R}^{d}$ that is invariant under the maximal compact subgroup $K$. We say that a subset $X$ of $G$ does not contract $F$ with respect to $|\cdot|$ if $F$ is invariant under all elements of $X$ and if $|g\cdot x|\geq|x|$ for all $g\in X$ and all $x\in F$. In what follows, we will simply say that $X$ does not contract $F$. To prove Proposition 3.1, we first establish the stronger result given by Proposition B, inspired by arguments from [Tho15]. To fix notation, let $p\in\llbracket 1,d\rrbracket$, and consider $\operatorname{Gr}_{p}(\mathbb{R}^{d})$, the Grassmannian of $p$-dimensional subspaces, note that the linear action of $\mathrm{SL}(d,\mathbb{R})$ induces a natural action on $\operatorname{Gr}_{p}(\mathbb{R}^{d})$. We say that $E,F\in\operatorname{Gr}_{p}(\mathbb{R}^{d})$ are transverse if $E+F=\mathbb{R}^{d}$.
The proof of Proposition B requires some preparation. Throughout this section, we work under the assumptions of Proposition B. Write $\Omega=\Gamma\cdot\mathcal{C}$ with $\mathcal{C}\subset\Omega$ compact. Let $\varepsilon>0$ be such that $\mathcal{C}+B_{\varepsilon}\subset\Omega$, where $B_{\varepsilon}$ denotes the Euclidean ball of radius $\varepsilon$ centered at 0, defined with respect to the Euclidean norm $\lvert\cdot\rvert$.
Lemma 3.2.
Let $y\in\partial\Omega$. Then there exist sequences $(g_{n})_{n\in\mathbb{N}}\subset\Gamma$, $(x_{n})_{n\in\mathbb{N}}\subset\mathcal{C}$, and $(H_{n})_{n\in\mathbb{N}}\subset\mathcal{F}$ such that, up to extraction of a subsequence:
  1. (1)
    $y_{n}:=g_{n}\cdot x_{n}$ converges to $y$ as $n\to\infty$;
  2. (2)
    for every $n$, one has $(H_{n}+g_{n}\cdot x_{n})\cap B(g_{n}\cdot x,\varepsilon)\subset\Omega$;
  3. (3)
    $H_{n}$ converges in $\operatorname{Gr}_{p}(\mathbb{R}^{n})$ to some element $H_{y}$ in the closure of $\mathcal{F}$, with $(H_{y}+y)\cap B(y,\varepsilon)\subset\partial\Omega$, where $B(y,\varepsilon)$ is Euclidean ball of radius $\varepsilon$ centered at $y$.
Proof.
Since $\Omega=\Gamma\cdot\mathcal{C}$, there exist sequences $(x_{n})_{n\in\mathbb{N}}\subset\mathcal{C}$ and $(g_{n})_{n\in\mathbb{N}}\subset\Gamma$ such that $g_{n}\cdot x_{n}\to y$. This proves the first point. For all $n\in\mathbb{N}$, we have
$g_{n}\cdot(x_{n}+B_{\varepsilon})=y_{n}+L(g_{n})\cdot B_{\varepsilon}\subset\Omega.$ (1)
For each $n$, write the Cartan decomposition $L(g_{n})=w_{n}a_{n}w_{n}^{\prime}$, with $w_{n},w_{n}^{\prime}\in K$ and $a_{n}\in A^{+}(L(\Gamma))$. Let $H$ be the $p$-subspace appearing in the hypothesis of Proposition B, and set $H_{n}:=w_{n}\cdot H$. We claim that $L(g_{n})\cdot B_{\varepsilon}$ contains $H_{n}\cap B_{\varepsilon}$. Indeed, let $v\in w_{n}\cdot H\cap B_{\varepsilon}$ and set $x=w_{n}^{-1}v\in H$. Since $K$ preserves the Euclidean norm $\lvert\cdot\rvert$ and $a_{n}$ does not contract $H$, we have $\lvert a_{n}^{-1}x\rvert\leq\lvert x\rvert=\lvert v\rvert\leq\varepsilon$. Therefore $u:=w_{n}^{\prime-1}(a_{n}^{-1}x)$ belongs to $B_{\varepsilon}$, and $L(g_{n})u=w_{n}a_{n}w_{n}^{\prime}\bigl(w_{n}^{\prime-1}a_{n}^{-1}x\bigr)=w_{n}x=v$. Hence $w_{n}\cdot H\cap B_{\varepsilon}\subset L(g_{n})\cdot B_{\varepsilon}$, which proves the claim.
It follows from (1) that $\Omega$ contains $y_{n}+(H_{n}\cap B_{\varepsilon})$, establishing the second point. By compactness of $\overline{\mathcal{F}}$, we may assume (up to extracting a subsequence) that $H_{n}$ converges in the Grassmannian to some $H_{y}\in\overline{\mathcal{F}}$. Then $\overline{\Omega}$ contains $y+(H_{y}\cap B_{\varepsilon})$, and since $y\in\partial\Omega$, the inclusion must in fact be
$(H_{y}\cap B_{\varepsilon})+y\subset\partial\Omega.$
This proves the third point and completes the proof. ∎
Remark 3.3.
In the previous lemma, we do not use the transversality of the elements of the closure of $\mathcal{F}$. In particular, the lemma asserts that each point of the boundary contains a small neighborhood of an affine subspace.
The next step is to show the uniqueness of the subspace $H_{y}$. Before doing so, we record the following elementary fact about transversality.
Lemma 3.4.
Let $E,F\subset\mathbb{R}^{d}$ be linear subspaces with $E+F=\mathbb{R}^{d}$. Let $E_{n}\to E$ in the Grassmannian and $y_{n}\to y$ in $\mathbb{R}^{d}$. Fix $\varepsilon>0$. Then for all sufficiently large $n$ the sets
$(E_{n}+y_{n})\cap B(y_{n},\varepsilon)\qquad\text{and}\qquad(F+y)\cap B(y,\varepsilon)$
have nonempty intersection.
Proof.
Up to translation, we may assume that $y=0$. The property $E+F=\mathbb{R}^{d}$ is open for $E$, so eventually we have $E_{n}+F=\mathbb{R}^{d}$. Hence $(E_{n}+y_{n})\cap F$ is a nonempty affine subspace $G_{n}$ of $\mathbb{R}^{d}$, converging to $E\cap F$. The Hausdorff limit of $G_{n}$ contains 0, and $y_{n}\rightarrow 0$, so $d(G_{n},y_{n})\rightarrow 0$ (where $d(\cdot,\cdot)$ is the euclidean distance). Since $G_{n}$ is closed and convex, there exists $b_{n}\in G_{n}$ such that $d(y_{n},b_{n})=d(y_{n},G_{n})\rightarrow 0$. Eventually $b_{n}$ is clearly both in $(E_{n}+y_{n})\cap B(y_{n},\varepsilon)$ and in $F\cap B(0,\varepsilon)$. ∎
We arrive at the following lemma.
Lemma 3.5.
Let $y\in\partial\Omega$, and let $H_{y}$ be an element of the closure of $\mathcal{F}$ as in Lemma 3.2. Then $H_{y}$ is the unique element of $\overline{\mathcal{F}}$ with the property that there exists a neighborhood $\mathcal{V}$ of $y$ in $\Omega$ such that $(y+H_{y})\cap\mathcal{V}\subset\partial\Omega$.
Proof.
Let $(y_{n})_{n\in\mathbb{N}}$, $(H_{n})_{n\in\mathbb{N}}$, and $H_{y}$ be as in Lemma 3.2. By shrinking $\varepsilon$ if necessary, we may assume that $B(y,\varepsilon)\subset\mathcal{V}$.
We proceed by contradiction and assume the existence of a subspace $H\in\overline{\mathcal{F}}$, distinct from $H_{y}$, that satisfies the hypothesis of the lemma. Then $H_{y}$ and $H$ are transverse. Since $y_{n}\to y$ and $H_{n}\to H_{y}$, and using Lemma 3.4, we deduce that, for $n$ large enough, the set $(H_{n}+y_{n})\cap B(y_{n},\varepsilon)$ intersects $(H+y)\cap B(y,\varepsilon)$.
However, by Lemma 3.2, we have $(H_{n}+y_{n})\cap B(y_{n},\varepsilon)\subset\Omega$, while $(H+y)\cap B(y,\varepsilon)\subset(H+y)\cap\mathcal{V}\subset\partial\Omega$, a contradiction. ∎
The next lemma goes a step further by showing that $\partial\Omega$ contains an affine subspace.
Lemma 3.6.
Let $y\in\partial\Omega$, and let $H_{y}$ be as in Proposition 3.2. Then $H_{y}+y\subset\partial\Omega$.
Proof.
Let $\mathcal{U}$ be the set of points $z\in H_{y}+y$ such that $(H_{y}+y)\cap B(z,\varepsilon)\subset\partial\Omega$. We will show that $\mathcal{U}$ is both open and closed (and nonempty) in $H_{y}+y$, so that $\mathcal{U}=H_{y}+y$, which will complete the proof.
Clearly, the set $\mathcal{U}$ is closed and contains $y$. We now show that $\mathcal{U}$ is also open. Let $z\in\mathcal{U}$ and $z^{\prime}\in(H_{y}+y)\cap B(z,\varepsilon/2)$. Then $(H_{y}+y)\cap B(z^{\prime},\varepsilon/2)\subset\partial\Omega$. Thus, by Lemma 3.5, we have $H_{y}=H_{z^{\prime}}$. By the definition of $H_{z^{\prime}}$, this implies $H_{y}\cap B(z^{\prime},\varepsilon)=H_{z^{\prime}}\cap B(z^{\prime},\varepsilon)\subset\partial\Omega,$ hence $z^{\prime}\in\mathcal{U}$. Therefore $\mathcal{U}$ is open, which completes the proof. ∎
We now have all the tools to prove Proposition B.
Proof of Proposition B.
First, we claim that two affine subspaces contained in $\partial\Omega$ whose linear parts lie in $\overline{\mathcal{F}}$ must be parallel. By Lemma 3.6, there exists at least one affine subspace contained in $\partial\Omega$ whose linear part belongs to $\overline{\mathcal{F}}$.
Assume, for contradiction, that there exist two affine subspaces $H$ and $H^{\prime}$ contained in $\partial\Omega$ with distinct linear parts in $\overline{\mathcal{F}}$. Then $H$ and $H^{\prime}$ are transverse, so they intersect at some point $y\in\partial\Omega$, which contradicts Lemma 3.5. Therefore $H$ and $H^{\prime}$ must have the same linear part; in other words $H$ and $H^{\prime}$ are parallel.
Next, let $H_{0}$ denote the common linear part of the hyperplanes contained in $\partial\Omega$. Since $\Gamma$ preserves $\Omega$, the group $L(\Gamma)$ preserves $H_{0}$. Let $x\in\Omega$ and suppose that $(H_{0}+x)$ is not contained in $\Omega$. Then there exists $y\in(x+H_{0})\cap\partial\Omega$, and by the previous argument $(y+H_{0})\subset\partial\Omega$. Hence $x\in\partial\Omega$, a contradiction. This completes the proof. ∎
3.3. Reduction in ${\operatorname{SO}}(2,2)$-case
In this subsection, we apply Proposition B to the geometric setting of primary interest, namely $({\operatorname{SO}}(2,2)\ltimes\mathbb{R}^{2,2},\mathbb{R}^{2,2})$-geometry. Throughout the paper, we denote by $\mathbb{R}^{2,2}$ the affine space $\mathbb{R}^{4}$ endowed with a bilinear form of signature $(2,2)$. The isometry group preserving this bilinear form is identified with ${\operatorname{O}}(2,2)\ltimes\mathbb{R}^{4}$; the subgroup preserving orientation is ${\operatorname{SO}}(2,2)\ltimes\mathbb{R}^{4}$, and we denote by ${\operatorname{SO}}_{0}(2,2)\ltimes\mathbb{R}^{4}$ its identity component. We begin with the following classical result about isotropic planes in $\mathbb{R}^{2,2}$.
Proposition 3.7.
The space of isotropic planes in $\mathbb{R}^{2,2}$ is the union of two connected components. Each connected component is a closed $\mathrm{SO}_{0}(2,2)$-orbit in $\mathrm{Gr}_{2}(\mathbb{R}^{4})$, and every pair of distinct planes in the same orbit is transverse.
Proof.
In this proof, we use the following model of $\mathbb{R}^{2,2}$: consider $\mathbb{R}^{2}\oplus\mathbb{R}^{2}$ endowed with the bilinear form $\langle\cdot,\cdot\rangle_{2,2}:=g\oplus(-g)$, where $g(\cdot,\cdot)$ is the standard Euclidean inner product on $\mathbb{R}^{2}$. It is straightforward to check that for each $A\in O(2)$, the graph of $A$, defined as $\mathrm{graph}(A):=\{(x,Ax):x\in\mathbb{R}^{2}\}$, is an isotropic plane in $\mathbb{R}^{2,2}$. We claim that the map defined by
$\mathcal{I}:O(2)\longrightarrow\{\text{totally isotropic\penalty 10000\ $2$-planes in }\mathbb{R}^{2,2}\},\qquad A\longmapsto\mathrm{graph}(A),$
is a homeomorphism. Continuity and injectivity are immediate. To prove surjectivity, let $W$ be an isotropic plane in $\mathbb{R}^{2,2}$, and consider the projections $\pi_{1},\pi_{2}:\mathbb{R}^{2}\oplus\mathbb{R}^{2}\longrightarrow\mathbb{R}^{2}$ onto the first and second factors, respectively. We claim that $\pi_{1}|_{W}$ is injective. Indeed, if $(v,0)\in W$, the isotropy condition implies
$0=\langle(v,0),(v,0)\rangle_{2,2}=g(v,v),$
so $v=0$. Since $\dim W=2=\dim\mathbb{R}^{2}$, it follows that $\pi_{1}|_{W}$ is a linear isomorphism $W\cong\mathbb{R}^{2}$. Define $A:\mathbb{R}^{2}\to\mathbb{R}^{2}$ by $A=\pi_{2}\circ(\pi_{1}|_{W})^{-1}.$ By construction, we have $W=\{(x,Ax):x\in\mathbb{R}^{2}\}$. Therefore, the space of isotropic planes has exactly two connected components, corresponding to the two connected components of $O(2)$.
We now show that each of these components is a single closed $\mathrm{SO}_{0}(2,2)$-orbit. First, since $\mathrm{SO}_{0}(2,2)$ is connected, every $\mathrm{SO}_{0}(2,2)$-orbit is connected; hence each orbit is contained in a single connected component. To show transitivity on each component, consider the subgroup, consider the subgroup
$H=\mathrm{SO}(2)\times\mathrm{SO}(2)\subset\mathrm{SO}_{0}(2,2)$
embedded diagonally via $(M,N)\mapsto\mathrm{diag}(M,N)$. Its action on graphs is given by
$(M,N)\cdot\mathrm{graph}(A)=\mathrm{graph}(NAM^{-1}).$
If $A,B\in O(2)$ lie in the same connected component of $O(2)$, then ${\operatorname{det}}B={\operatorname{det}}A$, so $BA^{-1}\in\mathrm{SO}(2)$. Taking $M=I$ and $N=BA^{-1}$ yields
$(I,BA^{-1})\cdot\mathrm{graph}(A)=\mathrm{graph}(B).$
Hence $H$, and therefore $\mathrm{SO}_{0}(2,2)$, acts transitively on each connected component. Consequently, each connected component of the space of isotropic planes is a single $\mathrm{SO}_{0}(2,2)$-orbit. Next, note that $O(2)$ is compact, so each of its components is compact, and hence its image under the homeomorphism $\mathcal{I}$ is compact and therefore closed in the Grassmannian. This shows that each orbit is closed.
Finally, we show that elements within each connected component are transverse. Consider the component corresponding to $\mathrm{SO}(2)$. Let $A,B\in\mathrm{SO}(2)$ with $A\neq B$. If $\mathrm{graph}(A)\cap\mathrm{graph}(B)\neq\{0\}$, then there exists a nonzero $v\in\mathbb{R}^{2}$ such that $Av=Bv$, i.e. $(AB^{-1})v=v$. But the only element of $\mathrm{SO}(2)$ having $1$ as an eigenvalue is the identity, hence $A=B$, a contradiction. Thus $\mathrm{graph}(A)$ and $\mathrm{graph}(B)$ are transverse. ∎
Proof of Proposition 3.1.
Let $\Omega$ and $\Gamma$ be as in Proposition 3.1. In this proof, we consider ${\operatorname{SO}}_{0}(2,2)$, the identity component of the linear group preserving the quadratic form of signature $(2,2)$ given by $q=dxdz+dydt$. For this quadratic form, each vector of the canonical basis $(e_{1},e_{2},e_{3},e_{4})$ of $\mathbb{R}^{4}$ is isotropic. Let us consider the Cartan subgroup of ${\operatorname{SO}}_{0}(2,2)$ given in the basis $(e_{1},e_{2},e_{3},e_{4})$ by
$A=\left\{\mathrm{diag}(\lambda,\mu,\lambda^{-1},\mu^{-1})\mid\lambda,\mu>0\right\},$
and the subset
$A^{+}=\left\{\mathrm{diag}(\lambda,\mu,\lambda^{-1},\mu^{-1})\mid\lambda,\mu\geq 1\right\}.$
Considering the maximal compact subgroup $K:=S(\operatorname{O}(2)\times\operatorname{O}(2))={\operatorname{SO}}_{0}(2,2)\cap{\operatorname{SO}}(4)$ of ${\operatorname{SO}}_{0}(2,2)$, we have a Cartan decomposition ${\operatorname{SO}}_{0}(2,2)=KA^{+}K$, in the sense of Section 3.1.
The set $A^{+}$ does not contract the isotropic plane $\mathrm{Span}(e_{1},e_{2})$ (with respect to the usual euclidean norm of $\mathbb{R}^{4}$). Let $\mathcal{F}$ be a closed ${\operatorname{SO}}_{0}(2,2)$-orbit of isotropic planes. Thus, we are in the context of Proposition B, and the proof follows. ∎
3.4. Stabilizers of an isotropic plane in $\mathbb{R}^{2,2}$
We collect here the Lie-theoretic computations used in the sequel. We consider on $\mathbb{R}^{4}$ the flat metric of signature $(2,2)$ given by
q=dxdz+dydt.q=dx\,dz+dy\,dt.
We also fix the isotropic plane $P_{0}=\mathbb{R}^{2}\times\{(0,0)\}$. Let $\mathcal{P}_{2}$ be the subgroup of ${\operatorname{SO}}_{0}(2,2)\ltimes\mathbb{R}^{2,2}$ whose linear part preserves $P_{0}$, and denote this linear part by $L(\mathcal{P}_{2})$.
Lemma 3.8.
[…]
Proof.
A straightforward block computation shows that any linear transformation in ${\operatorname{SO}}_{0}(2,2)$ preserving $P_{0}$ must be of the above block form for some $B\in{\operatorname{GL}}_{2}(\mathbb{R})$ and $b\in\mathbb{R}$. Since we work in the identity component ${\operatorname{SO}}_{0}(2,2)$, the matrix $B$ must have positive determinant. ∎
Remark 3.9.
Another convenient model for $\mathbb{R}^{2,2}$ is the vector space $\mathrm{M}_{2}(\mathbb{R})\cong\mathbb{R}^{4}$ equipped with the quadratic form given by the determinant. The group ${\operatorname{SL}}_{2}(\mathbb{R})\times{\operatorname{SL}}_{2}(\mathbb{R})$ acts on $\mathrm{M}_{2}(\mathbb{R})$ by left and right multiplication and preserves the determinant, yielding a surjective homomorphism
${\operatorname{SL}}_{2}(\mathbb{R})\times{\operatorname{SL}}_{2}(\mathbb{R})\longrightarrow{\operatorname{SO}}_{0}(2,2)$
with finite kernel. Under this identification, the subgroup $L(\mathcal{P}_{2})$ corresponds to a subgroup isomorphic to $B\times{\operatorname{SL}}_{2}(\mathbb{R})$, where $B$ is the upper-triangular Borel subgroup of ${\operatorname{SL}}_{2}(\mathbb{R})$.
We now investigate the nilradical of $\mathcal{P}_{2}$.
Definition 3.10.
We denote by $N$ the nilradical of $\mathcal{P}_{2}$. Explicitly, one has
$N=U\ltimes\mathbb{R}^{4},\quad\text{with}\quad U=\left\{\begin{pmatrix}I_{2}&bJ\\[4.0pt] 0&I_{2}\end{pmatrix}\;\Big|\;b\in\mathbb{R}\right\}.$
The group $N$ is a $5$-dimensional, index-22 nilpotent Lie group. Writing an element of $N$ as $(b,v_{1},v_{2})$ with $b\in\mathbb{R}$ and $v_{1},v_{2}\in\mathbb{R}^{2}$, one checks that
[…]
This leads to the following result.
Lemma 3.11.
[…]
$M\cdot(b,v_{1},v_{2})=(b\ {\operatorname{det}}M,\ Mv_{1},\ M^{-T}v_{2}).$ (3)
Proof.
Using (2), one checks that the map
$(M,(b,v_{1},v_{2}))\longmapsto\left(\begin{pmatrix}M&bMJ\\[5.69054pt] 0&M^{-T}\end{pmatrix},\ (v_{1},v_{2})\right)$
is a Lie group isomorphism from ${\operatorname{GL}}^{+}_{2}(\mathbb{R})\ltimes N$ onto $\mathcal{P}_{2}$. A direct computation shows that the multiplication and conjugation laws match. ∎
Having established this isomorphism, we now describe the Lie algebra $\mathfrak{n}$ of $N$. For $(s,v_{1},v_{2}),(t,w_{1},w_{2})\in\mathfrak{n}$, the bracket is
$(s,v_{1},v_{2}),(t,w_{1},w_{2})]=(0,\ sJw_{2}-tJv_{2},\ 0).$
Next, the infinitesimal action of $\mathfrak{gl}(2,\mathbb{R})$ on $\mathfrak{n}$ induced by (3) is
$X\cdot(s,v_{1},v_{2})=(s\,\operatorname{tr}X,\ Xv_{1},\ -X^{T}v_{2}).$
Thus the bracket in $\mathfrak{gl}(2,\mathbb{R})\ltimes\mathfrak{n}$ is
$(X,u),(Y,w)]=([X,Y],\ X\cdot w-Y\cdot u+[u,w]).$
Throughout the paper, we fix the basis of $\mathfrak{n}$ given by
$u$ $=(1,(0,0),(0,0)),$ $\qquad T_{1}$ $=(0,(1,0),(0,0)),$ $\qquad T_{2}$ $=(0,(0,1),(0,0)),$ (4)
$T_{3}$ $=(0,(0,0),(1,0)),$ $\qquad T_{4}$ $=(0,(0,0),(0,1)).$
The only nonzero brackets in this basis are
$u,T_{3}]=-T_{2},\qquad[u,T_{4}]=T_{1}.$ (5)
As consequence, we could show that $N$ is a homothety Lie group (see Definition 2.8).
Lemma 3.12.
The nilradical $N$ is a homothety Lie group. More precisely, for each $\lambda>1$ there exists an automorphism $\Phi_{\lambda}\in{\sf{Aut}}(N)$ whose differential at the identity is diagonalizable with eigenvalues of modulus >1>1, and which commutes with the ${\operatorname{GL}}_{2}^{+}(\mathbb{R})$-action on $N$.
Proof.
Define $\Psi_{\lambda}:\mathfrak{n}\to\mathfrak{n}$ by
$\Psi_{\lambda}(u)=\lambda u,\qquad\Psi_{\lambda}(T_{i})=\lambda^{2}T_{i}\ (i=1,2),\qquad\Psi_{\lambda}(T_{i})=\lambda T_{i}\ (i=3,4).$
One verifies that $\Psi_{\lambda}$ is an automorphism of the Lie algebra $\mathfrak{n}$ and that it commutes with $\mathrm{Ad}_{g}:\mathfrak{n}\to\mathfrak{n}$ for all $g\in{\operatorname{GL}}_{2}^{+}(\mathbb{R})$. Since $N$ is simply connected and nilpotent, $\Psi_{\lambda}$ integrates to a unique Lie group automorphism $\Phi_{\lambda}$, defined by $\Phi_{\lambda}(\exp X)=\exp(\Psi_{\lambda}(X))$, which moreover commutes with the ${\operatorname{GL}}_{2}^{+}(\mathbb{R})$-action on $N$. This completes the proof. ∎
Next, we consider the infinitesimal generator
$u_{1}=\begin{pmatrix}0&1\\ 0&0\end{pmatrix}\in\mathfrak{gl}(2,\mathbb{R}),$
corresponding to the nilpotent subgroup
[…]
Its nonzero brackets with the basis of $\mathfrak{n}$ are
$u_{1},T_{2}]=T_{1},\qquad[u_{1},T_{3}]=-T_{4}.$ (7)
In the end of this section, we introduce two natural projections that we use extensively in the proof of our theorem:
  • The projection $p:\mathcal{P}_{2}\to\mathcal{P}_{2}/N\cong{\operatorname{GL}}_{2}^{+}(\mathbb{R})$ modulo the nilradical.
  • The linear action of $\mathcal{P}_{2}$ on $\mathbb{R}^{2,2}$ induces an action on $\mathbb{R}^{2}\cong\mathbb{R}^{2,2}/P_{0}$. This action is given by the natural projection $q:\mathcal{P}_{2}\to{\operatorname{GL}}_{2}^{+}(\mathbb{R})\ltimes_{\mathsf{t}}\mathbb{R}^{2}$, where the action of ${\operatorname{GL}}_{2}^{+}(\mathbb{R})$ on $\mathbb{R}^{2}$ is defined by
    $M\cdot v=M^{-T}v.$
    The kernel of $q$ is
    $I:={\operatorname{Ker}}(q)=\left\{\begin{pmatrix}I_{2}&bJ\\[2.84526pt] 0&I_{2}\end{pmatrix}\;\Big|\;b\in\mathbb{R}\right\}\ltimes P_{0}=U\times P_{0}.$ (8)
Remark 3.13.
We record two useful observations.
  • The linear part of the kernel $I$ fixes the isotropic plane $P_{0}$ pointwise. Hence the semidirect product in (8) is in fact a direct product, that is, $I=U\times P_{0}$, and therefore $I$ is abelian.
  • The subgroup $I$ preserves every isotropic plane parallel to $P_{0}$. In particular, if $\Gamma$ is a discrete subgroup of $\mathcal{P}_{2}$ dividing a domain $\Omega\subset\mathbb{R}^{2,2}$ foliated by isotropic planes parallel to $P_{0}$, then $\operatorname{cd}(\Gamma\cap I)\leq 2$. This follows from the fact that $\Gamma\cap I$ acts properly and discontinuously on each leaf of the foliation. Since $I$ is abelian, it follows that $\Gamma\cap I$ is either trivial, or isomorphic to $\mathbb{Z}$ or $\mathbb{Z}^{2}$.
For the reader’s convenience, we list below the main notations and Lie brackets used throughout the paper.
Basis of $\mathfrak{n}$ $u,\,T_{1},\,T_{2},\,T_{3},\,T_{4}$ as in (4)
Nonzero brackets of the basis of $\mathfrak{n}$ $[u,T_{3}]=-T_{2},\qquad[u,T_{4}]=T_{1}$ as in (5)
Nonzero brackets with $u_{1}$ $[u_{1},T_{2}]=T_{1},\qquad[u_{1},T_{3}]=-T_{4}$ as in (7)
Isotropic plane $P_{0}=\mathbb{R}^{2}\times\{(0,0)\}\cong{\operatorname{Span}}(T_{1},T_{2})$ which is the center of the group NN
[…]
Kernels of projections ${\operatorname{Ker}}(p)=N$ and $I={\operatorname{Ker}}(q)=U\times P_{0}$
Table 1. Lie brackets and notations
4. Domains invariant under one-parameter subgroups
The goal of this section is to study domains of $\mathbb{R}^{2}$ that are invariant under certain one-parameter groups of ${\operatorname{GL}}_{2}(\mathbb{R})\ltimes\mathbb{R}^{2}$, where ${\operatorname{GL}}_{2}(\mathbb{R})$ acts by the usual linear action on $\mathbb{R}^{2}$. We are particularly interested in two families of one-parameter groups that play a central role in our arguments.
The first family consists of one-parameter groups of pure translations. The second family consists of one-parameter groups of $U_{1}\ltimes\mathbb{R}^{2}$ which are not contained in $U_{1}$. They are defined as follows: for each choice of parameters $a,b,c\in\mathbb{R}$, set
$J(a,b,c)=\left\{g_{t}:=\left(\begin{pmatrix}1&at\\[2.0pt] 0&1\end{pmatrix},\begin{pmatrix}\tfrac{1}{2}ab\,t^{2}+ct\\[4.0pt] bt\end{pmatrix}\right)\;\Big|\;t\in\mathbb{R}\right\}.$ (9)
We first study the simpler situation of one parameter group of translations.
Proposition 4.1.
Let $\mathcal{O}$ be a domain in the affine plane. Assume that $\mathcal{O}$ is invariant under a non-discrete subgroup $\mathcal{T}$ of translations. Then either $\mathcal{O}$ is the entire plane, or its boundary $\partial\mathcal{O}$ consists of one or two parallel affine lines.
Proof.
Consider $\overline{\mathcal{T}}$, the topological closure of $\mathcal{T}$ in the group of all pure translations. We claim that $\overline{\mathcal{T}}$ preserves $\mathcal{O}$. Indeed, since $\mathcal{T}$ preserves $\mathcal{O}$, it also preserves its complement $\mathcal{O}^{c}$, which is closed. The closure $\overline{\mathcal{T}}$ therefore preserves the closed set $\mathcal{O}^{c}$, and hence preserves $\mathcal{O}$ as well.
By hypothesis $\overline{\mathcal{T}}$ contains a one-parameter subgroup of pure translations; we denote it by $\mathcal{T}_{0}$. In particular, the set $\mathcal{O}$ is foliated by the parallel affine lines that are the orbits of the $\mathcal{T}_{0}$-action on the plane. We call this foliation $\mathcal{L}$.
Considering the quotient $\mathcal{O}/\mathcal{L}$, we obtain a connected open subset $J$ of $\mathbb{R}$. Up to composition with an affine automorphism of the plane, the possible cases for $J$ are $\mathbb{R}$, $(0,1)$, or $(0,\infty)$.
Since $\mathcal{O}$ is affinely isomorphic to the product $J\times\ell$, where $\ell$ is any $\mathcal{T}_{0}$-orbit (an affine line), this completes the proof of the lemma. ∎
In particular, we deduce the following corollary.
Corollary 4.2.
Let $\Omega\subset\mathbb{R}^{2,2}$ be a domain foliated by isotropic planes parallel to $P_{0}$. Suppose that $\Omega$ is divided by a discrete subgroup $\Gamma\leq{\operatorname{SO}}_{0}(2,2)\ltimes\mathbb{R}^{2,2}$, and that $\overline{q(\Gamma)}$ contains a one-parameter group of pure translations. Then $\Omega=\mathbb{R}^{2,2}$.
To prove this, let $\widehat{\Omega}$ denote the projection of $\Omega$ in $\mathbb{R}^{2,2}/P_{0}$ modulo the isotropic plane $P_{0}$. Then $\widehat{\Omega}$ is invariant under $q(\Gamma)$.
Proof.
Let $\mathcal{T}_{0}\leq q(\Gamma)$ be a one-parameter group of pure translations. Since $\widehat{\Omega}$ is $\mathcal{T}_{0}$-invariant, Lemma 4.1 implies that $\widehat{\Omega}$ is either the whole plane, in which case $\Omega=P_{0}\times\widehat{\Omega}=\mathbb{R}^{4}$, or that $\partial\widehat{\Omega}$ consists of one or two lines. As $\widehat{\Omega}$ is $q(\Gamma)$-invariant, it follows—after passing to a finite-index subgroup if necessary—that $q(\Gamma)$ preserves an affine line $l$. Consequently, $\Gamma$ preserves the hyperplane $\pi^{-1}(l)$ in $\mathbb{R}^{2,2}$, where $\pi:\mathbb{R}^{4}\to\mathbb{R}^{4}/P_{0}$ is the natural projection, contradicting Theorem 2.2. ∎
The following proposition deals with the subgroups $J(a,b,c)$ given in (9).
Proposition 4.3.
Let $a,b,c\in\mathbb{R}$ with $b\neq 0$, and let $\mathcal{O}\subset\mathbb{R}^{2}$ be a domain invariant under $J(a,b,c)$. Then $\mathcal{O}$ is contractible.
Proof.
Let $g_{t}$ be an element of $J(a,b,c)$ (see (9)). The linear part of $g_{t}$ acts on $\mathbb{R}^{2}$ by the inverse transpose, which gives a map
$g_{t}:(x,y)\longmapsto(\,x+at\,y+\frac{1}{2}ab\,t^{2}+ct,\;y+bt\,).$
Let $\pi_{2}:\mathbb{R}^{2}\to\mathbb{R}$ be the projection on the second coordinate. For any $(x,y)\in\mathbb{R}^{2}$ and $t\in\mathbb{R}$ we have $\pi_{2}(g_{t}(x,y))=y+bt.$ Since $b\neq 0$ and $\mathcal{O}$ is nonempty and invariant, the set $\pi_{2}(\mathcal{O})$ is nonempty and invariant under all translations by $bt$. Hence $\pi_{2}(\mathcal{O})=\mathbb{R}$. In particular, there exists $(0,s_{0})\in\mathcal{O}$. Set $L_{0}=\{(x,0):y\in\mathbb{R}\}$ and $S:=\mathcal{O}\cap L_{0}$. Then $S$ is a nonempty open subset of $L_{0}$. Define
$\Phi:\mathbb{R}\times S\longrightarrow\mathcal{O},\qquad\Phi(t,(s,0)):=g_{t}(s,0).$
We claim that $\Phi$ is a homeomorphism. It is clearly continuous, and it has inverse
$\Phi^{-1}(x,y)=\left(\frac{y}{b},\,g_{-y/b}(x,y)\right),$
which is continuous. Hence $\Phi$ is a homeomorphism, so $\mathcal{O}\cong\mathbb{R}\times S$. Since $\mathcal{O}$ is connected, the set $S$ must be connected, hence an open interval $I\subset L_{0}$. Therefore $\mathcal{O}\cong\mathbb{R}\times I,$ which is contractible. This proves the proposition. ∎
Remark 4.4.
It is worth noting that one parameter groups of $U_{1}\ltimes_{\mathsf{t}}\mathbb{R}^{2}$, which are not contained in $U_{1}$ are of the form
$J^{\mathsf{t}}(a,b,c)=\left\{\left(\begin{pmatrix}1&at\\[2.0pt] 0&1\end{pmatrix},\begin{pmatrix}bt\\[4.0pt] -\tfrac{1}{2}ab\,t^{2}+ct\end{pmatrix}\right)\;\Big|\;t\in\mathbb{R}\right\}.$
In particular, the conclusion of Proposition 4.3 still holds. Namely any domain $\mathcal{O}$ invariant under $J^{\mathsf{t}}(a,b,c)$ is contractible.
We record the following corollary.
Corollary 4.5.
Let $\Omega\subset\mathbb{R}^{2,2}$ be a domain foliated by isotropic planes parallel to $P_{0}$. Suppose that $\Omega$ is divided by a discrete subgroup $\Gamma\leq{\operatorname{SO}}_{0}(2,2)\ltimes\mathbb{R}^{2,2}$, and that $\overline{q(\Gamma)}$ contains a one parameter group of the from $J^{\mathsf{t}}(a,b,c)$ with $b\neq 0$. If $\Gamma$ is solvable then $\Omega=\mathbb{R}^{2,2}$.
Proof.
Since $J^{\mathsf{t}}(a,b,c)$ preserves $\hat{\Omega}$, Proposition 4.3 and Remark 4.4 imply that $\hat{\Omega}$ is contractible. Since $\Omega$ is foliated by planes parallel to $P_{0}$, we have $\Omega\cong P_{0}\times\hat{\Omega}$, hence $\Omega$ is contractible. Therefore $\operatorname{cd}(\Gamma)=4$. Finally, the solvability of $\Gamma$, together with Theorem 2.5, implies completeness. ∎
5. The quotient geometry: Non-injective projection
The goal of this section is to prove the following result.
Proposition 5.1.
Let $\Omega\subset\mathbb{R}^{2,2}$ be a domain foliated by isotropic planes parallel to $P_{0}$. Suppose that $\Omega$ is divided by a discrete subgroup $\Gamma\leq{\operatorname{SO}}_{0}(2,2)\ltimes\mathbb{R}^{2,2}$, and that the restriction $q:\Gamma\to{\operatorname{GL}}^{+}_{2}(\mathbb{R})\ltimes_{\mathsf{t}}\mathbb{R}^{2}$ is not injective. Then $\Omega=\mathbb{R}^{2,2}$.
Note that by hypothesis we have either $\Gamma\cap I\cong\mathbb{Z}$ or $\Gamma\cap I\cong\mathbb{Z}^{2}$ (recall Remark 3.13). The proof of Proposition 5.1 will proceed by distinguishing these two possibilities for $\Gamma\cap I$.
5.1. Rank two kernel
In this part we prove the following:
Proposition 5.2.
Let $\Gamma$ and $\Omega$ be as in Proposition 5.1. Assume moreover that $\Gamma\cap I\cong\mathbb{Z}^{2}$. Then $\Omega=\mathbb{R}^{2,2}$.
The proof follows from the following general fact.
Lemma 5.3.
Let $G$ and $H$ be Lie groups acting on manifolds $X$ and $Y$ respectively. Let $\Gamma$ be a discrete subgroup of $G$, and let $\rho:\Gamma\to H$ be a group homomorphism. Let $F:X\to Y$ be a fibration which is $\Gamma$-equivariant in the sense that $F(\gamma\cdot x)=\rho(\gamma)\cdot F(x)$ for all $\gamma\in\Gamma$ and $x\in X$. Assume that:
  1. (1)
    $\Gamma$ acts properly discontinuously and cocompactly on $X$;
  2. (2)
    ${\operatorname{Ker}}(\rho)$ preserves each fiber of $F$ and acts properly discontinuously and cocompactly on it.
Then $\rho(\Gamma)$ is discrete and the induced action of $\rho(\Gamma)$ on $Y$ is properly discontinuous.
Proof.
Let $C\subset Y$ be compact and set $A:=F^{-1}(C)\subset X$. By hypothesis, each fiber of $F$ modulo $\Gamma_{0}$ is compact, and therefore $A/\Gamma_{0}$ is compact. Hence there exists a compact set $B\subset A$ such that $\Gamma_{0}\cdot B=A$. Define $S:=\{\sigma\in\Gamma:\ \sigma\cdot B\cap B\neq\varnothing\}$. Because $B$ is compact and $\Gamma$ acts properly discontinuously on $X$, the set $S$ is finite.
We claim that
$\{h\in\rho(\Gamma)\mid h\cdot C\cap C\neq\varnothing\}\subset\rho(S).$
Let $h\in\rho(\Gamma)$ satisfy $h\cdot C\cap C\neq\varnothing$. Choose $\gamma\in\Gamma$ with $\rho(\gamma)=h$. Then there exist $y,y^{\prime}\in C$ with $h\cdot y=y^{\prime}$. Pick $x\in F^{-1}(y)\subset A$. Then $F(\gamma\cdot x)=\rho(\gamma)\cdot F(x)=h\cdot y=y^{\prime}$, so $\gamma\cdot x\in A$. Now let $k_{1},k_{2}\in\Gamma_{0}$ and $b_{1},b_{2}\in B$ such that $x=k_{1}\cdot b_{1}$ and $\gamma\cdot x=k_{2}b_{2}$. Then $(k_{2}^{-1}\gamma k_{1})\cdot b_{1}=b_{2}$. Therefore $k_{2}^{-1}\gamma k_{1}\in S$ and so $\gamma\in k_{2}Sk_{1}^{-1}$. Because $k_{1},k_{2}\in{\operatorname{Ker}}\rho$, we have $\rho(\gamma)\in\rho(S)$. Hence $h=\rho(\gamma)\in\rho(S)$, proving the claim. Since $\rho(S)$ is finite, the set $\{h\in\rho(\Gamma)\mid h\cdot C\cap C\neq\varnothing\}$ is finite, and so the action of $\rho(\Gamma)$ on $Y$ is properly discontinuous. The discreteness follows in the same way, indeed if $\rho(\gamma_{n})\to{\operatorname{Id}}$, then for any $y\in Y$ $\rho(\gamma_{n})\cdot y\to y$, take a compact neighborhood $C^{\prime}$ of $y$, then $\rho(\gamma_{n})\cdot y\in C^{\prime}$ for large $n$, hence for large $n$, $\rho(\gamma_{n})\in\{h\in\rho(\Gamma)\mid h\cdot C^{\prime}\cap C^{\prime}\neq\varnothing\}$ which is finite set, and so $\rho(\gamma_{n})={\operatorname{Id}}$ for $n$ large enough. ∎
Proof of Proposition 5.2.
By Lemma 5.3, the discrete group $q(\Gamma)$ acts properly discontinuously and cocompactly on $\widehat{\Omega}\subset\mathbb{R}^{2}$. This implies that $\widehat{\Omega}$ is either the affine plane $\mathbb{R}^{2}$, a half-plane $H$, a quarter-plane $Q$, or the once-punctured plane $\mathbb{R}^{2}\setminus\{0\}$ (see, for instance, [Bau14, Proposition 5.2] or [Ben60, §9.1]). If $\widehat{\Omega}$ is not the whole plane, then the boundary of $\Omega$, which is given by $P_{0}\times\partial\widehat{\Omega}$, is a proper algebraic set preserved by $\Gamma$, contradicting Theorem 2.2. Hence $\widehat{\Omega}=\mathbb{R}^{2}$, and thus $\Omega=\mathbb{R}^{4}$. This completes the proof. ∎
5.2. Rank one kernel
In this subsection, we prove Proposition 5.1 under the assumption that $\Gamma\cap I\cong\mathbb{Z}$.
Proposition 5.4.
Let $\Gamma$ and $\Omega$ be as in Proposition 5.1. Assume moreover that $\Gamma\cap I\cong\mathbb{Z}$. Then $\Omega=\mathbb{R}^{2,2}$.
The proof of the proposition requires some preparation. We start by this lemma.
Lemma 5.5.
Let $P_{0}={\operatorname{Span}}(T_{1},T_{2})$ be the center of $N$ and assume that $\Gamma\cap I\cong\mathbb{Z}$. Then
  • If $\Gamma\cap P_{0}$ is trivial, then after passing to a subgroup of index $2$ of $\Gamma$, we have $p(\Gamma)\leq{\operatorname{SL}}(2,\mathbb{R})$.
  • If $\Gamma\cap P_{0}$ is nontrivial, then after passing to a subgroup of index $2$ of $\Gamma$, we may assume that $p(\Gamma)$ lies in
    $B_{1}:=\left\{\begin{pmatrix}1&x\\ 0&\lambda\end{pmatrix}\;\middle|\;\lambda>0,\,x\in\mathbb{R}\right\}.$ (10)
Proof.
Consider $S={\operatorname{Span}}(\Gamma\cap I)\leq I$, which is the Malcev closure of $\Gamma\cap I$ in $I$, and denote by $\mathfrak{s}$ its Lie algebra. Since $\Gamma$ normalizes $\Gamma\cap I\cong\mathbb{Z}$, it also normalizes $S$. In particular, this yields a homomorphism $\Gamma\to{\sf{Aut}}(\mathbb{Z})\cong\{\pm 1\}$. The kernel of this homomorphism is a subgroup of $\Gamma$ of index at most $2$ and acts trivially on $\Gamma\cap I$. Therefore, it induces the trivial action on $S$. On the level of Lie algebras, this gives $\mathrm{Ad}_{\gamma}(u)=u$ for all $u\in\mathfrak{s}$ and all $\gamma\in\Gamma$.
Write $\gamma=(B,n)\in\Gamma$, and let $u_{0}=(s_{0},v_{0})$ be a generator of $\mathfrak{s}$. A direct computation gives
$\mathrm{Ad}_{\gamma}(u_{0})=\left({\operatorname{det}}(B)\,s_{0},\;B\cdot(v_{0}-s_{0}Jw_{2})\right).$
The equation $\mathrm{Ad}_{\gamma}(u_{0})=u_{0}$ therefore implies
${\operatorname{det}}(B)\,s_{0}=s_{0},\quad B\bigl(v_{0}-s_{0}Jw_{2}\bigr)=v_{0},\quad\forall\,B\in p(\Gamma).$ (11)
If $\Gamma\cap P_{0}$ is trivial, then $s_{0}\neq 0$, and hence ${\operatorname{det}}(B)=1$ for all $B\in p(\Gamma)$, so $p(\Gamma)\leq{\operatorname{SL}}(2,\mathbb{R})$. If $\Gamma\cap P_{0}$ is nontrivial, then $s_{0}=0$, and equation (11) reduces to $Bv_{0}=v_{0}$, with $v_{0}\in P_{0}$. Up to conjugacy, we may assume that $v_{0}=T_{1}$, and hence $p(\Gamma)$ is contained in the group $B_{1}$. ∎
Having established this, we proceed with the proof of Proposition 5.4 by considering two cases, depending on whether $\Gamma\cap P_{0}$ is trivial or not.
5.2.1. $\Gamma\cap P_{0}$ trivial
We start with the following result.
Proposition 5.6.
If $\Gamma\cap I\cong\mathbb{Z}$ and $\Gamma\cap P_{0}$ is trivial, then $\Omega=\mathbb{R}^{2,2}$.
From now on, and until the end of the proof of Proposition 5.6, we work under the assumptions above. We define $L:{\operatorname{GL}}_{2}^{+}(\mathbb{R})\ltimes N\to{\operatorname{GL}}_{2}^{+}(\mathbb{R})\ltimes U$ so that $L(\Gamma)$ is the linear part of $\Gamma$. Since $\Gamma\cap P_{0}$ is trivial, Lemma 5.5 implies that $p(\Gamma)\leq{\operatorname{SL}}_{2}(\mathbb{R})$, and thus $L(\Gamma)\leq{\operatorname{SL}}_{2}(\mathbb{R})\ltimes U$ (see 3.10 for the definition of $U$). However, elements of ${\operatorname{SL}}_{2}(\mathbb{R})$ commute with $U$, and therefore $L(\Gamma)$ is contained in the direct product ${\operatorname{SL}}_{2}(\mathbb{R})\times U$. Next, we define
$\Gamma_{0}=\Gamma\cap\mathbb{R}^{4},$ (12)
the subgroup of pure translations in $\Gamma$. This subgroup is invariant under the action of $L(\Gamma)$. Moreover, it is a discrete subgroup of $\mathbb{R}^{4}$, and hence $\operatorname{cd}(\Gamma_{0})\leq 4$.
Lemma 5.7.
If $\operatorname{cd}(\Gamma_{0})\geq 3$, then $\Omega=\mathbb{R}^{2,2}$.
Proof.
We consider the natural projection $\mathbb{R}^{4}\to{\operatorname{Span}}(T_{3},T_{4})$, with kernel $P_{0}={\operatorname{Span}}(T_{1},T_{2})$. By hypothesis, $\Gamma\cap P_{0}$ is trivial. Hence, we obtain an injective homomorphism $\Gamma_{0}\to{\operatorname{Span}}(T_{3},T_{4})$. Since $\operatorname{cd}(\Gamma_{0})\geq 3$, the image of $\Gamma_{0}$ cannot be discrete in ${\operatorname{Span}}(T_{3},T_{4})$. Therefore, $\overline{q(\Gamma_{0})}$ contains a one-parameter group of pure translations, and completeness then follows from Corollary 4.2. ∎
Next we investigate the rank two case.
Lemma 5.8.
If $\operatorname{cd}(\Gamma_{0})=2$, then $\Omega=\mathbb{R}^{2,2}$.
For that, we need the following lemma about invariant planes under $U$.
Sublemma 5.8.1.
Let $M\in U$ be different from the identity. Then the invariant planes of $M$ are
$P_{0}={\operatorname{Span}}(T_{1},T_{2})\quad\text{or}\quad{\operatorname{Span}}((Jv,0),(u,v)),$
where $v\neq 0$ and $u\in\mathbb{R}^{2}$.
Proof.
Let $b\neq 0$, and consider $M$, the non trivial element of $U$ given by
$M=\begin{pmatrix}I_{2}&bJ\\ 0&I_{2}\end{pmatrix}\in U,$
Let $(u,v)\in\mathbb{R}^{4}$, then
M(u,v)(u,v)=(bJv,0)M(u,v)-(u,v)=(bJv,0) (13)
Let $V\subset\mathbb{R}^{4}$ be a 22–dimensional linear subspace invariant under $M$. If $\pi_{2}(V)$ is trivial, then $V\subset\mathbb{R}^{2}\oplus\{0\}$. Since $\dim V=2$, this forces $V=\mathbb{R}^{2}\oplus\{0\}=P_{0}$. If not, choose $(u,v)\in V$ with $v\neq 0$. By (13), $(Jv,0)\in V$ and hence the vectors $(Jv,0)$ and $(u,v)$ are linearly independent in $V$. Since $\dim V=2$, it follows that $V={\operatorname{Span}}\{(Jv,0),(u,v)\}.$
Proof of Lemma 5.8.
Consider a nontrivial element $\gamma\in\Gamma\cap I$. Since $I=U\times P_{0}$, the linear part of $\gamma$ is a nontrivial element of $U$, which preserves $\Gamma_{0}$ and hence preserves the plane $S={\operatorname{Span}}(\Gamma_{0})$. Since $\dim(S)=2$, by Sublemma 5.8.1 we have either $S={\operatorname{Span}}(T_{1},T_{2})=P_{0}$ or $S={\operatorname{Span}}((Jv,0),(u,v))$. The first case is excluded since $\Gamma\cap P_{0}$ is trivial.
For the second case, we have $q(\Gamma_{0})\leq q(S)\leq{\operatorname{Span}}(v)\cong\mathbb{R}$. Since $\Gamma_{0}\cap I=\Gamma_{0}\cap P_{0}$ is trivial, we have $\Gamma_{0}\cong q(\Gamma_{0})$, and hence $q(\Gamma_{0})$ cannot be discrete; otherwise we would obtain a discrete subgroup of $\mathbb{R}$ with cohomological dimension $2$, which is impossible. Completeness now follows from Corollary 4.2. ∎
The next lemma excludes the case $\operatorname{cd}(\Gamma_{0})=1$.
Lemma 5.9.
$\operatorname{cd}(\Gamma_{0})$ cannot be equal to $1$,
Proof.
Consider a nontrivial element $\gamma\in\Gamma\cap I$. Since $I=U\times P_{0}$, the linear part of $\gamma$ is a nontrivial element of $U$, which preserves $\Gamma_{0}$ and hence preserves the line $S={\operatorname{Span}}(\Gamma_{0})$. However, it is not difficult to check that any line invariant under a nontrivial element of $U$ must lie in the plane $P_{0}$. Hence $\Gamma\cap I\subset P_{0}$, which is a contradiction. ∎
We now turn to the case where $\Gamma_{0}$ is trivial.
Proposition 5.10.
If $\Gamma_{0}$ is trivial, then $\Omega=\mathbb{R}^{2,2}$.
Throughout this paper, we define
$T:=q(\Gamma)\cap\bigl(\{\mathrm{Id}\}\times\mathbb{R}^{2}\bigr)=q(\Gamma\cap N).$ (14)
The subgroup of pure translations in $q(\Gamma)$. We begin by establishing the following result.
Lemma 5.11.
If $\Gamma_{0}$ is trivial, then either $\operatorname{cd}(\Gamma\cap N)=1$ or $T$ is not discrete.
Proof.
Assume that $T$ is a discrete subgroup of $\mathbb{R}^{2}$. Then $\operatorname{cd}(T)\leq 2$. Using the short exact sequence
$1\to\Gamma\cap I\to\Gamma\cap N\to T\to 1,$
we obtain $\operatorname{cd}(\Gamma\cap N)\leq 3$. Since $\Gamma\cap I\leq\Gamma\cap N$, we also have $\operatorname{cd}(\Gamma\cap N)\geq 1$. In what follows, denote by $S$ the Malcev closure of $\Gamma\cap N$ in $N$. Observe that $\dim(S)=\operatorname{cd}(\Gamma\cap N)$. Moreover, $\Gamma\cap N$ is abelian, since
$\Gamma\cap N,\Gamma\cap N]\leq\Gamma_{0},$
which is trivial by assumption. Therefore, $\Gamma\cap N$ is a uniform lattice in the simply connected nilpotent Lie group $S$, which implies that $S$ itself is abelian.
  • If $\operatorname{cd}(\Gamma\cap N)=3$, then $\dim(S)=3$. It follows from Lemma 8.2 that the Lie algebra of $S$ must be ${\mathfrak{a}}_{2}={\operatorname{Span}}(u+t,\;T_{1},\;T_{2}),$ with $t\in{\operatorname{Span}}(T_{3},T_{4})$. This implies
    $T\leq q(S)\leq{\operatorname{Span}}(t),$
    in particular $\operatorname{cd}(T)\leq 1$. This is impossible, as $\operatorname{cd}(\Gamma\cap N)=3$ and $\operatorname{cd}(\Gamma\cap N)\leq 1+\operatorname{cd}(T)\leq 2,$ a contradiction.
  • If $\operatorname{cd}(\Gamma\cap N)=2$, then $\dim(S)=2$. It follows from Lemma 8.3 that the Lie algebra of $S$ must be equal to ${\operatorname{Span}}(u+t,\;t^{\prime}),$ with $t\in{\operatorname{Span}}(T_{3},T_{4})$ and $t^{\prime}\in{\operatorname{Span}}(T_{1},T_{2})$. We consider the connected subgroup $I\cap S$ of $N$. This subgroup is nontrivial, as it contains $\Gamma\cap I$. Hence its dimension is either $1$ or $2$. If its dimension is $1$, then necessarily $I\cap S={\operatorname{Span}}(t^{\prime})$, and therefore $\Gamma\cap I\leq{\operatorname{Span}}(t^{\prime})$. This is a contradiction, since $\Gamma\cap P_{0}$ is trivial.
    Thus $I\cap S$ has dimension $2$ and hence must coincide with $S$. This implies that $\Gamma\cap I=\Gamma\cap N$. Indeed, we always have $\Gamma\cap I\leq\Gamma\cap N$. For the reverse inclusion, since $S=I\cap S\leq I$, we obtain $\Gamma\cap N\leq\Gamma\cap S\leq\Gamma\cap I$. But this is a contradiction, as $\operatorname{cd}(\Gamma\cap N)=2$ and $\operatorname{cd}(\Gamma\cap I)=1$. This completes the proof.
The next result shows that $p(\Gamma)$ cannot be discrete if $\operatorname{cd}(\Gamma\cap N)=1$.
Lemma 5.12.
If $\operatorname{cd}(\Gamma\cap N)=1$, then $p(\Gamma)$ is not discrete.
Proof.
Assume by contradiction that $p(\Gamma)$ is discrete. By Lemma 5.5, $p(\Gamma)$ is a discrete subgroup of ${\operatorname{SL}}_{2}(\mathbb{R})$ and hence up to finite index, $p(\Gamma)$ acts properly and freely on the contractible space $\mathbb{H}^{2}$, this implies that $\operatorname{cd}(p(\Gamma))\leq 2$. Using the short exact sequence
$1\to\Gamma\cap N\to\Gamma\to p(\Gamma)\to 1,$
we deduce that $\operatorname{cd}(\Gamma)\leq 3$, which contradicts Theorem 2.3. ∎
Lemma 5.13.
If $\operatorname{cd}(\Gamma\cap N)=1$, then $\overline{p(\Gamma)}^{\circ}$ is a one parameter group of ${\operatorname{SL}}_{2}(\mathbb{R})$.
Proof.
By Lemma 3.12, we may apply Theorem C to the group ${\operatorname{GL}}_{2}^{+}(\mathbb{R})\ltimes N$, to deduce that $\overline{p(\Gamma)}^{\circ}$ is a nilpotent subgroup of ${\operatorname{SL}}_{2}(\mathbb{R})$. By Lemma 5.12, it is not trivial and hence its dimension is one; otherwise, it would not be nilpotent. This completes the proof. ∎
We start by considering the simpler case where $\overline{p(\Gamma)}^{\circ}$ is not a parabolic subgroup of ${\operatorname{SL}}_{2}(\mathbb{R})$.
Lemma 5.14.
If $\overline{p(\Gamma)}^{\circ}$ is not a one-parameter parabolic subgroup of ${\operatorname{SL}}_{2}(\mathbb{R})$, then $\Omega=\mathbb{R}^{2,2}$.
Proof.
Since $\overline{p(\Gamma)}^{\circ}$ is a one parameter group of ${\operatorname{SL}}_{2}(\mathbb{R})$, which is either elliptic or hyperbolic, then its normalizers is abelian. Since $\overline{p(\Gamma)}$ normalizes its identity component, it follows that $\overline{p(\Gamma)}$ is abelian. As consequence, $L(\Gamma)\leq p(\Gamma)\times U$ is abelian, and since $\Gamma_{0}$ is trivial, we conclude that $\Gamma\cong L(\Gamma)$ is abelian. The completeness follows from Theorem 2.4. ∎
Next, we deal with the parabolic case. Up to conjugacy, we may assume that $\overline{p(\Gamma)}^{\circ}=U_{1}$ (see (6)).
Proposition 5.15.
If $\overline{p(\Gamma)}^{\circ}=U_{1}$, then $\Omega=\mathbb{R}^{2,2}$.
First, note that the normalizers of $U_{1}$ is given by
$B=\left\{\begin{pmatrix}e^{t}&x\\ 0&e^{-t}\end{pmatrix}\ \Big|\ t,x\in\mathbb{R}\right\},$
and hence $\overline{p(\Gamma)}\leq B$. We write $B$ as the semidirect product $A\ltimes\mathbb{R}$ where the $A$-factor corresponds to the hyperbolic one-parameter group given by
$A=\{\mathrm{diag}(e^{t},e^{-t})\mid t\in\mathbb{R}\}.$
We show the following.
Lemma 5.16.
If $\overline{p(\Gamma)}^{\circ}=U_{1}$, then either $\overline{p(\Gamma)}$ is nilpotent or $\overline{p(\Gamma)}\cong\mathbb{Z}\ltimes U_{1}$.
Proof.
Consider the natural projection $B\to A$. Since the kernel of this projection is $U_{1}$, which equals $\overline{p(\Gamma)}^{\circ}$, the projection of $\overline{p(\Gamma)}$ is discrete in $A\cong\mathbb{R}$. If the projection to $A$ is trivial then $\overline{p(\Gamma)}$ is nilpotent. If not the projection is isomorphic to $\mathbb{Z}$, and hence $\overline{p(\Gamma)}\cong\mathbb{Z}\ltimes U_{1}.$
Proof of Proposition 5.15.
Since $\Gamma_{0}$ is trivial, we have $\Gamma\cong L(\Gamma)$. This implies that $\Gamma$ is solvable, as $L(\Gamma)\leq B\times U$ is solvable. If $\operatorname{cd}(\Gamma)=4$ then completeness follows from Theorem 2.5. Assume now that $\operatorname{cd}(\Gamma)>4$, and consider $\Gamma_{1}=\Gamma\cap(U_{1}\ltimes N)$. From the short exact sequence
$1\to\Gamma_{1}\to\Gamma\to\mathbb{Z}\to 1,$
we deduce that $\operatorname{cd}(\Gamma_{1})\geq 4$. Moreover, observe that $\Gamma_{1},\Gamma_{1}]\leq\Gamma_{1}\cap\mathbb{R}^{4}\leq\Gamma_{0}$, which is trivial by assumption; hence $\Gamma_{1}$ is abelian. Let $S_{1}$ be the syndetic hull of $\Gamma_{1}$. Then $S_{1}$ is abelian; this follows from the fact that $S_{1}$ is a simply connected nilpotent group containing an abelian uniform lattice.
Next, since $\dim(S_{1})=\operatorname{cd}(\Gamma_{1})=4$, it follows from Lemma 8.1 that $S_{1}=\mathbb{R}^{4}$, and so $\Gamma_{1}\leq\mathbb{R}^{4}$. This contradicts the fact that $\Gamma_{0}$ is trivial. ∎
Proof of Proposition 5.10.
By Lemma 5.11, we have either $\operatorname{cd}(\Gamma\cap N)=1$ or $T$ is not discrete. If $T$ is not discrete, completeness follows from Corollary 4.2. If $\operatorname{cd}(\Gamma\cap N)=1$, then $\overline{p(\Gamma)}^{\circ}$ is a one-parameter subgroup of ${\operatorname{SL}}_{2}(\mathbb{R})$. If it is not parabolic, completeness follows from Lemma 5.14. If it is parabolic, completeness follows from Proposition 5.15. This completes the proof. ∎
We can finally prove Proposition 5.6.
Proof of Proposition 5.6.
The proof follows by combining Lemmas 5.7, 5.9, and 5.8 in the case where the translation group $\Gamma_{0}$ is nontrivial. The conclusion in the case where $\Gamma_{0}$ is trivial follows from Proposition 5.10. ∎
5.2.2. $\Gamma\cap P_{0}$ is nontrivial
In this part we will show this result.
Proposition 5.17.
If $\Gamma\cap I\cong\mathbb{Z}$ and $\Gamma\cap P_{0}$ is nontrivial, then $\Omega=\mathbb{R}^{2,2}$.
By Lemma 5.5, we may assume that $p(\Gamma)\leq B_{1}$. Note that $B_{1}$ is isomorphic to the affine group and, in particular, it is not unimodular. It may be written as the semidirect product $\mathbb{R}^{+}\ltimes U_{1}$, where $U_{1}$ is the unipotent subgroup defined in (6). From now on, and until the end of the proof of Proposition 5.17, we work under the assumptions that $\Gamma\cap I\cong\mathbb{Z}$ and $p(\Gamma)\leq B_{1}$.
The first step is to prove the following.
Proposition 5.18.
If either $p(\Gamma)$ or $q(\Gamma)$ is discrete, then $\Omega=\mathbb{R}^{2,2}$.
We start with the case where $p(\Gamma)$ is discrete.
Lemma 5.19.
If $p(\Gamma)$ is discrete, then $\Omega=\mathbb{R}^{2,2}$.
Proof.
We claim that if $p(\Gamma)$ is discrete, then either $T$ is non-discrete or $T\cong\mathbb{Z}^{2}$ (recall the definition of $T$ in (14)). First note that we may assume that $p(\Gamma)$ is torsion-free, up to passing to a finite-index subgroup of $\Gamma$. Using the short exact sequence $1\to\Gamma\cap N\to\Gamma\to p(\Gamma)\to 1$, we obtain $\mathrm{cd}(\Gamma)\leq\mathrm{cd}(\Gamma\cap N)+\mathrm{cd}(p(\Gamma))$.
By contradiction, assume that $T$ is discrete with $\operatorname{cd}(T)\leq 1$, then we have the short exact sequence $1\to\Gamma\cap I\to\Gamma\cap N\to T\to 1$. Thus $\mathrm{cd}(\Gamma\cap N)\leq\mathrm{cd}(\Gamma\cap I)+\mathrm{cd}(T)\leq 2.$ Since $p(\Gamma)$ is a discrete subgroup of $B_{1}$, we must have $\mathrm{cd}(p(\Gamma))=1$. Indeed, if $\mathrm{cd}(p(\Gamma))=2$, then $p(\Gamma)$ would be a uniform lattice in $B_{1}$, which is impossible because $B_{1}$ is not unimodular. Hence $\mathrm{cd}(\Gamma)\leq 3,$ contradicting Theorem 2.3. This proves the claim.
If $T$ is non-discrete, then completeness follows from Corollary 4.2. If $T$ is discrete, then we must have $T\cong\mathbb{Z}^{2}$. Now, observe that the linear part of $q(\Gamma)\leq B_{1}\ltimes_{\mathsf{t}}\mathbb{R}^{2}$, which is by definition $p(\Gamma)$, preserves the group of pure translations $T\cong\mathbb{Z}^{2}$. Hence $p(\Gamma)$ is volume-preserving (i.e., has determinant one), which implies that $p(\Gamma)\leq U_{1}$. Consequently, $\Gamma\leq U_{1}\ltimes N$ is nilpotent, and completeness follows from Theorem 2.4. ∎
We now consider the case where $q(\Gamma)$ is discrete.
Lemma 5.20.
If $q(\Gamma)$ is discrete, then $\Omega=\mathbb{R}^{2,2}$.
The proof uses the following fact about $B_{1}\ltimes_{\mathsf{t}}\mathbb{R}^{2}$.
Sublemma 5.20.1.
The group $B_{1}\ltimes_{\mathsf{t}}\mathbb{R}^{2}$ is not unimodular.
Proof.
Fix $\lambda\neq 1$ and consider the diagonal matrix $M=\mathrm{diag}(1,\lambda)\in B_{1}$. We will show that the determinant of $\mathrm{Ad}_{M}$ is not equal to $1$. For that, consider the basis
$u_{1}=\begin{pmatrix}0&1\\[4.0pt] 0&0\end{pmatrix},\qquad u_{2}=\begin{pmatrix}0&0\\[4.0pt] 0&1\end{pmatrix}$
of the Lie algebra of $B_{1}$. We compute $\mathrm{Ad}_{M}$ in the basis $(u_{1},u_{2},T_{3},T_{4})$. A simple computation gives $Mu_{1}M^{-1}=(1/\lambda)u_{1}$ and $Mu_{2}M^{-1}=u_{2}$. Therefore the adjoint action $\mathrm{Ad}_{M}$ in the basis $(u_{1},u_{2},T_{3},T_{4})$ is
$\mathrm{Ad}_{M}=\begin{pmatrix}\frac{1}{\lambda}&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&\frac{1}{\lambda}\end{pmatrix}$.
Thus ${\operatorname{det}}(\mathrm{Ad}_{M})=\frac{1}{\lambda^{2}}\neq 1$, which completes the proof. ∎
Proof of Lemma 5.20.
We first claim that if $q(\Gamma)$ is discrete, then $\mathrm{cd}(q(\Gamma))\leq 3$. Since up to finite index, $q(\Gamma)$ acts properly on the contractible Lie group $B_{1}\ltimes_{\mathsf{t}}\mathbb{R}^{2}$ of dimension $4$, we always have $\mathrm{cd}(q(\Gamma))\leq 4$. Suppose for contradiction that $\mathrm{cd}(q(\Gamma))=4$. Then $q(\Gamma)$ would be a uniform lattice in $B_{1}\ltimes_{\mathsf{t}}\mathbb{R}^{2}$, contradicting non-unimodularity (Sublemma 5.20.1), hence $\mathrm{cd}(q(\Gamma))=3$. The short exact sequence
$1\longrightarrow\Gamma\cap I\longrightarrow\Gamma\longrightarrow q(\Gamma)\longrightarrow 1$
then implies $\mathrm{cd}(\Gamma)\leq\mathrm{cd}(q(\Gamma))+\mathrm{cd}(\Gamma\cap I)\leq 3+1=4$. In the other hand $\Gamma\leq B_{1}\ltimes N$ is solvable. Thus $\Omega=\mathbb{R}^{2,2}$ by Theorem 2.5. ∎
We can now conclude the proof of Proposition 5.18.
Proof of Proposition 5.18.
The proof follows by combining Lemma 5.19 and Lemma 5.20. ∎
The rest of this section is devoted to proving the following.
Proposition 5.21.
If both $q(\Gamma)$ and $p(\Gamma)$ are non-discrete, then $\Omega=\mathbb{R}^{2,2}$.
We start by showing this result.
Lemma 5.22.
If $p(\Gamma)$ is not discrete, then either $\overline{p(\Gamma)}=U_{1}$ or $\overline{p(\Gamma)}\cong\mathbb{Z}\ltimes U_{1}$.
The proof of Lemma 5.22 proceeds by showing that the identity component $\overline{p(\Gamma)}^{\circ}$ coincides with the nilradical of $B_{1}$.
Sublemma 5.22.1.
The group $\overline{p(\Gamma)}^{\circ}$ coincides $U_{1}$
Proof.
We consider the discrete group $\Gamma_{nd}=\Gamma\cap\left(\overline{p(\Gamma)}^{\circ}\ltimes N\right)$. By Lemma 3.12, we may apply Theorem C to the group $\overline{p(\Gamma)}^{\circ}\ltimes N$ to get a nilpotent syndetic hull $S$ in $\overline{p(\Gamma)}^{\circ}\ltimes N$ with the property that $\overline{p(S)}=\overline{p(\Gamma)}^{\circ}$
Now, by contradiction, assume that $\overline{p(\Gamma)}^{\circ}$ is different from $U_{1}$. Then it is either equal to $B_{1}$ or to a one-parameter subgroup of $B_{1}$ not contained in $U_{1}$. The case $\overline{p(\Gamma)}^{\circ}=B_{1}$ is impossible, since $B_{1}$ is not nilpotent whereas $\overline{p(S)}=\overline{p(\Gamma)}^{\circ}$ is nilpotent.
Assume now that $\overline{p(\Gamma)}^{\circ}$ is a one-parameter subgroup not contained in the nilradical of $B_{1}$. Then the normalizer of $\overline{p(\Gamma)}^{\circ}$ is itself. However $\overline{p(\Gamma)}$ normalizes its identity component; hence $\overline{p(\Gamma)}\leq\overline{p(\Gamma)}^{\circ}$. This implies that $\overline{p(\Gamma)}$ is connected, so that $\overline{p(\Gamma)}=\overline{p(\Gamma)}^{\circ}$ and therefore $\Gamma_{nd}=\Gamma$. Moreover, $p(S)$ is closed, being a connected one-dimensional Lie group, and hence $p(S)=\overline{p(\Gamma)}$. By Theorem 2.3, $\operatorname{cd}(\Gamma)\geq 4$ and so $\dim S\geq\operatorname{cd}(\Gamma_{nd})\geq 4$. Since $p(S)$ is one-dimensional, we deduce that $\dim(S\cap N)\geq 3$. Now we rewrite $B_{1}\ltimes N$ is its original form, namely $(B_{1}\ltimes U)\ltimes\mathbb{R}^{4}$, see Definition 3.10. The nilpotency of $S$ implies that the $L(S)$-action on $S\cap\mathbb{R}^{4}$, is unipotent, where $L$ is the projection into the linear part. Now, let $\gamma=(M,v)\in L(\Gamma)\leq L(S)$ such that $M$ is non-unipotent element of $B_{1}$ and $v\in U$. Then the adjoint action of $\gamma$ in the basis $T_{1},T_{2},T_{3},T_{4}$ of $\mathbb{R}^{4}$ (see Table 1) is given by
$\mathrm{Ad}_{\gamma}=\begin{pmatrix}1&\star&\star&\star\\ 0&\lambda&\star&0\\ 0&0&1&0\\ 0&0&\star&\frac{1}{\lambda}\end{pmatrix}$.
We conclude that the maximal dimension of a subgroup of $\mathbb{R}^{4}$ on which the $\mathrm{Ad}_{\gamma}$–action is unipotent is two, given by ${\operatorname{Span}}\{T_{1},T_{3}\}$. Hence, $S\cap\mathbb{R}^{4}$ has dimension exactly two, and in particular $\dim(S\cap N)=3$. This implies that $S\cap U$ is nontrivial. For any $g\in S\cap U$ with $\neq\mathrm{Id}, gg$ preserves the plane $S\cap\mathbb{R}^{4}={\operatorname{Span}}\{T_{1},T_{3}\}$. This contradicts Sublemma 5.8.1. ∎
We can now prove Lemma 5.22.
Proof of Lemma 5.22.
Consider the natural projection $B_{1}\to\mathbb{R}^{+}$. Since the kernel of this projection is $U_{1}$, which equals to $\overline{p(\Gamma)}^{\circ}$, the projection of $\overline{p(\Gamma)}$ is discrete in $\mathbb{R}^{+}$. If the projection to $\mathbb{R}^{+}$ is trivial then $\overline{p(\Gamma)}=U_{1}$. If not the projection is isomorphic to $\mathbb{Z}$, and hence $\overline{p(\Gamma)}\cong\mathbb{Z}\ltimes U_{1}.$
We now have all the tools to prove Proposition 5.21.
Proof of Proposition 5.21.
By assumption $p(\Gamma)$ is not discrete, so by Lemma 5.22, either $\overline{p(\Gamma)}=U_{1}$ is nilpotent or $p(\Gamma)\cong\mathbb{Z}\ltimes U_{1}.$ If $\overline{p(\Gamma)}=U_{1}$, then $\Gamma\leq U_{1}\ltimes N$ is nilpotent. Completeness then follows from Theorem 2.4. Now consider the case $p(\Gamma)\cong\mathbb{Z}\ltimes U_{1}$ and let
$\Lambda=q(\Gamma)\cap(U_{1}\ltimes_{\mathsf{t}}\mathbb{R}^{2}),\quad\text{and}\quad H=\overline{\Lambda}^{\circ}$
be its identity component. In particular, we have $\overline{q(\Gamma)}^{\circ}=H.$ By assumption $q(\Gamma)$ is not discrete, so $H$ is a non-trivial connected closed connected subgroup of $U_{1}\ltimes_{\mathsf{t}}\mathbb{R}^{2}$. We proceed according to the dimension of $H$:
  • If $\dim(H)\geq 2$, then $H\cap\mathbb{R}^{2}\neq\{0\}$, in particular $\overline{q(\Gamma)}$ contains a one-parameter subgroup of pure translations, and completeness follows by Corollary 4.2.
  • If $\dim(H)=1$, then $H$ is a one-parameter subgroup of $U_{1}\ltimes_{t}\mathbb{R}^{2}$. We distinguish two cases:
    1. (1)
      If $H$ is a one-parameter family of pure translations, completeness follows from Corollary 4.2.
    2. (2)
      Otherwise, one has $H=\overline{q(\Gamma)}^{\circ}=J^{\mathsf{t}}(a,b,c)$ for some $a\neq 0$ and $b,c\in\mathbb{R}$, see Remark 4.4. We claim that $b\neq 0$.
      Suppose by contradiction that $b=0$. One checks that the normalizer of $J^{\mathsf{t}}(a,0,c)$ in $U_{1}\ltimes_{\mathsf{t}}\mathbb{R}^{2}$ is $U_{1}\ltimes_{\mathsf{t}}{\operatorname{Span}}(T_{4})$. Since $\overline{\Lambda}$ normalizes its identity component, we have $\overline{\Lambda}\leq U_{1}\ltimes_{\mathsf{t}}{\operatorname{Span}}(T_{4})$, and therefore $q(\Gamma)\leq\mathbb{Z}\ltimes(U_{1}\ltimes_{\mathsf{t}}{\operatorname{Span}}(T_{4})).$ Let $\widehat{\gamma}\in q(\Gamma)$ be an element projecting to a generator of the $\mathbb{Z}$-factor. Up to conjugation by an element of $U_{1}\ltimes_{\mathsf{t}}\mathbb{R}^{2}$, we may assume that
      $\widehat{\gamma}=\left(\begin{pmatrix}1&0\\ 0&\lambda\end{pmatrix},\begin{pmatrix}\alpha\\ 0\end{pmatrix}\right)$,
      where $\alpha\in\mathbb{R}$. Assume first that $\alpha=0$. Since conjugation by an element of $U_{1}\ltimes_{\mathsf{t}}\mathbb{R}^{2}$ preserves $U_{1}\ltimes_{\mathsf{t}}{\operatorname{Span}}(T_{4})$, it follows that all elements of $q(\Gamma)$ are contained in
      $\left\{\begin{pmatrix}1&\star\\ 0&\lambda^{n}\end{pmatrix}\;\Big|\;n\in\mathbb{Z}\right\}\ltimes_{\mathsf{t}}{\operatorname{Span}}(T_{4}).$
      Consequently $\Gamma$ is contained in the preimage under $q$ of $(\mathbb{Z}\ltimes U_{1})\ltimes_{\mathsf{t}}{\operatorname{Span}}(T_{4})$, namely,
      $\left\{\begin{pmatrix}1&\star&\star&\star\\ 0&\lambda^{n}&\star&\star\\ 0&0&1&0\\ 0&0&\star&\frac{1}{\lambda^{n}}\end{pmatrix}\mid n\in\mathbb{Z}\right\}\ltimes\begin{pmatrix}\star\\ \star\\ 0\\ \star\end{pmatrix}$.
      This group preserves the affine plane $\{x_{3}=0\}$, and hence $\Gamma$ preserves this plane. This contradicts Theorem 2.2.
      We now assume that $\alpha\neq 0$. Since $\Lambda=q(\Gamma)\cap(U_{1}\ltimes_{\mathsf{t}}\mathbb{R}^{2})$ is normal in $q(\Gamma)$, it is in particular invariant under $\mathrm{Ad}_{\widehat{\gamma}}$. Its closure is therefore also $\mathrm{Ad}_{\widehat{\gamma}}-$invariant, and hence its identity component $H$ is invariant as well. Expressing the adjoint action of $\widehat{\gamma}$ on the invariant plane spanned by $U_{1}$ and $T_{4}$, we obtain
      $\begin{pmatrix}\frac{1}{\lambda}&0\\ \frac{\alpha}{\lambda}&\frac{1}{\lambda}\end{pmatrix}$.
      The only invariant direction of this matrix is $T_{4}$. Thus $H=J^{\mathsf{t}}(a,b,c)$ must be tangent to $T_{4}$, which forces $a=0$, a contradiction.
Having established this, Proposition 5.4 then follows by combining Propositions 5.18 and 5.21. We can now complete the proof of the main result of the section.
Proof of Proposition 5.1.
Let $\Gamma$ and $\Omega$ be as in Proposition 5.1. Then either $\Gamma\cap I$ is isomorphic to $\mathbb{Z}$ or to $\mathbb{Z}^{2}$. In both cases, completeness follows from Propositions 5.2 and 5.4. ∎
6. The quotient geometry: Injective projection
In this section, we consider the situation opposite to that of Proposition 5.1. Our goal is to prove the following result.
Proposition 6.1.
Let $\Omega\subset\mathbb{R}^{2,2}$ be a domain foliated by isotropic planes parallel to $P_{0}$. Suppose that $\Omega$ is divided by a discrete subgroup $\Gamma\leq{\operatorname{SO}}_{0}(2,2)\ltimes\mathbb{R}^{2,2}$, and that the restriction $q:\Gamma\to{\operatorname{GL}}^{+}_{2}(\mathbb{R})\ltimes_{\mathsf{t}}\mathbb{R}^{2}$ is injective. Then $\Omega=\mathbb{R}^{2,2}$.
We now proceed to the proof of Proposition 6.1, distinguishing cases according to the structure of the group of pure translations $T$ in $q(\Gamma)$ defined in (14). Recall that the linear part of $q(\Gamma)\leq{\operatorname{GL}}_{2}^{+}(\mathbb{R})\ltimes_{\mathsf{t}}\mathbb{R}^{2}$, namely $p(\Gamma)$, preserves the subgroup $T$. When $T$ is not discrete, completeness follows from Corollary 4.2. Thus, it suffices to focus on the case where $T\leq\mathbb{R}^{2}$ is discrete. In this case, either $T$ is trivial, or $T\cong\mathbb{Z}$, or $T\cong\mathbb{Z}^{2}$. We start by excluding the $\mathbb{Z}^{2}$ case.
Lemma 6.2.
If $\Gamma\cap I$ is trivial, then we cannot have $T\cong\mathbb{Z}^{2}$.
Proof.
Assume by contradiction that $T\cong\mathbb{Z}^{2}$. Since $T$ is $p(\Gamma)$-invariant, it follows that, up to conjugacy, we have $p(\Gamma)\leq{\operatorname{SL}}_{2}(\mathbb{Z})$. By the injectivity of $q$, we have $q(\Gamma)\cong\Gamma$, and thus there is a well-defined projection $q(\Gamma)\to p(\Gamma)$ with kernel $T$. This gives the short exact sequence:
$1\to T\cong\mathbb{Z}^{2}\to q(\Gamma)\cong\Gamma\to p(\Gamma)\to 1.$
Since $p(\Gamma)\leq{\operatorname{SL}}_{2}(\mathbb{Z})$, then up to finite index, $\operatorname{cd}(p(\Gamma))\leq 1$ and so $\mathrm{cd}\,\Gamma\leq 3$. This contradicts Theorem 2.3. ∎
6.0.1. Rank one translation group
The goal of this section is to prove the following.
Proposition 6.3.
If $T\cong\mathbb{Z}$ and $\Gamma\cap I$ is trivial, then $\Omega=\mathbb{R}^{2,2}$.
The same analysis as in Subsection 5.2 shows that, after possibly passing to a subgroup of index at most $2$, we may assume up to conjugacy that $T={\operatorname{Span}}(T_{3})$, and $p(\Gamma)$ fixes $T_{3}$. In particular $p(\Gamma)$ is contained in the group
$B_{1}^{*}:=\left\{\begin{pmatrix}1&0\\ x&\lambda\end{pmatrix}\;\Bigg|\;\lambda\in\mathbb{R}_{>0},\ x\in\mathbb{R}\right\}\cong\mathbb{R}^{+}\ltimes U_{1}^{*},$ (15)
where $U_{1}^{*}$ denotes the nilradical of $B_{1}^{*}$. From now on, and until the end of the proof of Proposition 6.3, we work under the assumptions that $\Gamma\cap I$ is trivial and $T\cong\mathbb{Z}$.
Lemma 6.4.
The groups $p(\Gamma)$ and $q(\Gamma)$ are not discrete.
Proof.
Assume by contradiction that $p(\Gamma)$ is discrete. From the short exact sequence $1\to\Gamma\cap N\to\Gamma\to p(\Gamma)\to 1,$ we obtain $\operatorname{cd}(\Gamma)\leq\operatorname{cd}(\Gamma\cap N)+\operatorname{cd}(p(\Gamma)).$ Since $\Gamma\cap I$ is trivial, we have $\Gamma\cap N\cong T$, and hence $\operatorname{cd}(\Gamma\cap N)=1$. Because $p(\Gamma)$ is a discrete subgroup of $B_{1}^{*}$, it follows that $\operatorname{cd}(p(\Gamma))=1$, and therefore $\operatorname{cd}(\Gamma)\leq 2,$ contradicting Theorem 2.3. We now show that $q(\Gamma)$ cannot be discrete. Assume again by contradiction that $q(\Gamma)$ is discrete. One can show, exactly as in the proof of Sublemma 5.20.1, that the group $B_{1}^{*}\ltimes_{\mathsf{t}}\mathbb{R}^{2}$ is not unimodular. Thus, as in the proof of Lemma 5.20, we deduce that $\operatorname{cd}(q(\Gamma))\leq 3$. But $\Gamma\cong q(\Gamma)$, so $\operatorname{cd}(\Gamma)=\operatorname{cd}(q(\Gamma))\leq 3$, once more contradicting Theorem 2.3. ∎
Next, we derive the following result.
Lemma 6.5.
We have either $\overline{p(\Gamma)}=U_{1}^{*}$ or $\overline{p(\Gamma)}\cong\mathbb{Z}\ltimes U_{1}^{*}$.
Proof.
Since $p(\Gamma)\leq B_{1}^{*}$ and $p(\Gamma)$ is not discrete, the same analysis as in Sublemma 5.22.1 shows that $\overline{p(\Gamma)}^{\circ}=U_{1}^{*}$. Consider the natural projection $B_{1}^{*}\to\mathbb{R}^{+}$, where recall that $B_{1}^{*}=\mathbb{R}^{+}\ltimes U_{1}^{*}$. Its kernel is $U_{1}^{*}$, which coincides with $\overline{p(\Gamma)}^{\circ}$. Hence, the projection of $\overline{p(\Gamma)}$ is discrete in $\mathbb{R}^{+}$. If this projection is trivial, then $\overline{p(\Gamma)}=U_{1}^{*}$; otherwise, $\overline{p(\Gamma)}\cong\mathbb{Z}\ltimes U_{1}^{*}$. ∎
Proof of Proposition 6.3.
If $\overline{p(\Gamma)}=U_{1}^{*}$, then $\Gamma\leq U_{1}^{*}\ltimes N$ is nilpotent and so the completeness follows from Theorem 2.4. Otherwise, by Lemmas 6.4 and 6.5, we have $\overline{p(\Gamma)}\cong\mathbb{Z}\ltimes U_{1}^{*}$ and $q(\Gamma)$ is not discrete. Let us consider the group
$\Lambda:=q(\Gamma)\cap(U_{1}^{*}\ltimes_{\mathsf{t}}\mathbb{R}^{2}),$
and let $H:=\overline{\Lambda}^{\circ}$ be its identity component. In particular, we have $\overline{q(\Gamma)}^{\circ}=H$. Since $q(\Gamma)$ is not discrete, then $H$ is a nontrivial connected closed subgroup of $U_{1}^{*}\ltimes_{\mathsf{t}}\mathbb{R}^{2}$. We proceed exactly as in the proof of Proposition 5.21.
  • If $\dim(H)\geq 2$, then $H\cap\mathbb{R}^{2}\neq\{0\}$. In particular $\overline{q(\Gamma)}$ contains a one-parameter group of pure translations, and completeness follows from Corollary 4.2.
  • If $\dim(H)=1$, then $H$ is a one-parameter subgroup of $U_{1}^{*}\ltimes_{\mathsf{t}}\mathbb{R}^{2}$. We distinguish two cases:
    1. (1)
      If $H$ is a one-parameter family of pure translations, completeness follows from Corollary 4.2.
    2. (2)
      Otherwise,
      $H=J_{*}^{\mathsf{t}}(a,b,c)=\left\{\left(\begin{pmatrix}1&0\\[2.0pt] at&1\end{pmatrix},\begin{pmatrix}-\tfrac{1}{2}ab\,t^{2}+ct\\[4.0pt] bt\end{pmatrix}\right)\;\middle|\;t\in\mathbb{R}\right\}$
      for some $a\neq 0$ and $b,c\in\mathbb{R}$. We claim that $b\neq 0$.
      Suppose by contradiction that $b=0$. One checks that the normalizer of $J_{*}^{\mathsf{t}}(a,0,c)$ in $U_{1}^{*}\ltimes_{\mathsf{t}}{\operatorname{Span}}(T_{3})$ is equal to $U_{1}^{*}\ltimes_{\mathsf{t}}{\operatorname{Span}}(T_{3})$. Since $\overline{\Lambda}$ normalizes its identity component, we obtain $\overline{\Lambda}\leq U_{1}^{*}\ltimes_{\mathsf{t}}{\operatorname{Span}}(T_{3}),$ and therefore $q(\Gamma)\leq\mathbb{Z}\ltimes(U_{1}^{*}\ltimes_{\mathsf{t}}{\operatorname{Span}}(T_{3}))$.
      Let $\widehat{\gamma}\in q(\Gamma)$ be an element projecting to a generator of the $\mathbb{Z}$-factor. Up to conjugacy by an element of $U_{1}^{*}\ltimes_{\mathsf{t}}{\operatorname{Span}}(T_{3})$, we may assume that
      $\widehat{\gamma}=\left(\begin{pmatrix}1&0\\ 0&\lambda\end{pmatrix},\begin{pmatrix}\alpha\\ 0\end{pmatrix}\right)$,
      where $\alpha\in\mathbb{R}$. Since conjugation by an element of $U_{1}^{*}\ltimes_{\mathsf{t}}\mathbb{R}^{2}$ preserves $U_{1}^{*}\ltimes_{\mathsf{t}}{\operatorname{Span}}(T_{3})$, it follows that $q(\Gamma)$ is contained in
      $\left\{\begin{pmatrix}1&0\\ \star&\lambda^{n}\end{pmatrix}\mid n\in\mathbb{Z}\right\}\ltimes_{\mathsf{t}}\left\{\begin{pmatrix}\star\\ 0\end{pmatrix}\right\}$,
      This implies that $\Gamma$ is contained in the preimage under $q$ of $(\mathbb{Z}\ltimes U_{1}^{*})\ltimes_{\mathsf{t}}{\operatorname{Span}}(T_{3})$, namely
      $\left\{\begin{pmatrix}1&0&\star&\star\\ \star&\lambda^{n}&\star&\star\\ 0&0&1&\star\\ 0&0&0&\frac{1}{\lambda^{n}}\end{pmatrix}\mid n\in\mathbb{Z}\right\}\ltimes\begin{pmatrix}\star\\ \star\\ \star\\ 0\end{pmatrix}$.
      This group preserves the affine plane $x_{4}=0$, so $\Gamma$ preserves that plane, contradicting Theorem 2.2. Therefore $b\neq 0$, and completeness follows from Corollary 4.5.
6.0.2. Trivial translation group
The next step is to understand what happens if $q$ is injective and the pure translation group $T$ is trivial. A first observation is that under these assumptions, the restriction of $p$ to $\Gamma$ is injective, that is $\Gamma\cap N$ is trivial, see Table 1 for the notation.
Proposition 6.6.
If $\Gamma\cap I$ and $T$ are trivial, then $\Omega=\mathbb{R}^{2,2}$.
From now on, and until the end of the proof of Proposition 6.6, we work under the assumption that both $\Gamma\cap I$ and $T$ are trivial. In this case, the restrictions of $p$ and $q$ to $\Gamma$ are injective. The first step is to show that $\Gamma$ is solvable.
Lemma 6.7.
The group $\Gamma$ is solvable.
We begin with the following observation.
Sublemma 6.7.1.
The group $p(\Gamma)$ cannot be discrete.
Proof.
By contradiction, if $p(\Gamma)$ is discrete, then $\operatorname{cd}\Gamma=\operatorname{cd}p(\Gamma)\leq 3$. The last inequality is due to the fact that, up to finite index, the group $p(\Gamma)$ acts properly and freely on the contractible space ${\operatorname{GL}}_{2}^{+}(\mathbb{R})/{\operatorname{SO}}(2)\cong\mathbb{R}^{+}\times\mathbb{H}^{2}$. ∎
Having established this, we now consider the group $\overline{p(\Gamma)}^{\circ}$, which is non-trivial since $p(\Gamma)$ is not discrete.
Sublemma 6.7.2.
The identity component $\overline{p(\Gamma)}^{\circ}$ cannot be equal to the group of pure homotheties in ${\operatorname{GL}}_{2}^{+}(\mathbb{R})$.
Proof.
We denote by $\mathbb{R}$ the subgroup of pure homotheties in ${\operatorname{GL}}_{2}^{+}(\mathbb{R})$. Assume for contradiction that $\overline{p(\Gamma)}^{\circ}=\mathbb{R}$. Let $\pi:{\operatorname{GL}}_{2}(\mathbb{R})\to{\operatorname{SL}}_{2}(\mathbb{R})$ be the natural projection. Since ${\operatorname{Ker}}(\pi)=\mathbb{R}=\overline{p(\Gamma)}^{\circ}$, it follows that $\pi(\overline{p(\Gamma)})$ is a discrete subgroup of ${\operatorname{SL}}_{2}(\mathbb{R})$. In particular $\pi(p(\Gamma))$ is also discrete and therefore, up to finite index, one has $\operatorname{cd}\big(\pi(p(\Gamma))\big)\leq 2.$ Consider $\Gamma^{\prime}=\Gamma\cap(\mathbb{R}\ltimes N).$ We claim that $\operatorname{cd}(\Gamma^{\prime})\geq 2.$ Indeed, from the short exact sequence
$1\longrightarrow\Gamma^{\prime}\longrightarrow\Gamma\longrightarrow\pi(p(\Gamma))\longrightarrow 1,$
we obtain $\operatorname{cd}(\Gamma)\leq\operatorname{cd}(\Gamma^{\prime})+\operatorname{cd}\big(\pi(p(\Gamma))\big).$ On the other hand, by Theorem 2.3, we know that $\operatorname{cd}(\Gamma)\geq 4.$ Thus $\operatorname{cd}(\Gamma^{\prime})\geq 2.$
Next, since $\Gamma^{\prime}=\Gamma\cap\bigl(\overline{p(\Gamma)}^{\circ}\ltimes N\bigr)$, Lemma 3.12 allows us to apply Theorem C to the group $\mathbb{R}\ltimes N$, and we obtain a nilpotent syndetic hull $S_{\Gamma^{\prime}}$ inside $\mathbb{R}\ltimes N$.
Choose $\gamma\in\Gamma^{\prime}$ such that the linear part of $\gamma$ is $L(\gamma)=(\widehat{\gamma},x)\in\mathbb{R}\ltimes U$, where $\widehat{\gamma}=\mathrm{diag}(\lambda,\lambda)$ with $\lambda\neq 1$. The adjoint action of $L(\gamma)$ on $\mathbb{R}^{4}$ with respect to the basis $T_{1},T_{2},T_{3},T_{4}$ (see Table 1) is given by
$\mathrm{Ad}_{L(\gamma)}=\begin{pmatrix}\lambda&\star&\star&\star\\ 0&\lambda&\star&0\\ 0&0&\frac{1}{\lambda}&0\\ 0&0&\star&\frac{1}{\lambda}\end{pmatrix}$. (16)
The subspace of fixed vectors of $\mathrm{Ad}_{L(\gamma)}$ is trivial. Since $S_{\Gamma^{\prime}}$ is nilpotent, the linear action of $L(S_{\Gamma^{\prime}})$ on $S_{\Gamma^{\prime}}\cap\mathbb{R}^{4}$ is unipotent. Applying this to $L(\gamma)\in L(S_{\Gamma^{\prime}})$ and using (16), we deduce that $S_{\Gamma^{\prime}}\cap\mathbb{R}^{4}$ is trivial. In particular, $S_{\Gamma^{\prime}}$ injects into $\mathbb{R}\ltimes U_{1}^{*}$. However, since $\dim S_{\Gamma^{\prime}}\geq 2$, it follows that $S_{\Gamma^{\prime}}\cong\mathbb{R}\ltimes U_{1}^{*}=B_{1}^{*}$. This is a contradiction, because $S_{\Gamma^{\prime}}$ is a unimodular Lie group, whereas $B_{1}^{*}$ is not. ∎
Proof of Lemma 6.7.
It is not difficult to check that the normalizer of a solvable subgroup of ${\operatorname{GL}}_{2}^{+}(\mathbb{R})$ is solvable, unless it is the group of pure homotheties. By Theorem C, the group $\overline{p(\Gamma)}^{\circ}$ is nilpotent and hence solvable. Since $\overline{p(\Gamma)}$ normalizes its identity component, the result then follows from Sublemma 6.7.2. ∎
By Theorem 2.3, we have $\operatorname{cd}(\Gamma)\geq 4$. Since $p$ is injective, it follows that $\Gamma\cong p(\Gamma)$, and hence $\Gamma$ is solvable. Completeness follows from Theorem 2.5 in the case $\operatorname{cd}(\Gamma)=4$. Completeness also holds when $p(\Gamma)$ is abelian, without any assumption on the cohomological dimension. Indeed, in this case $\Gamma\cong p(\Gamma)$ is abelian, and the conclusion follows from Theorem 2.4. The remaining and more delicate part is to establish the following.
Proposition 6.8.
If $p(\Gamma)$ is not abelian, then $\operatorname{cd}(\Gamma)=4$.
The remainder of this section is devoted to proving the proposition and so from now on we assume that $p(\Gamma)$ is not abelian. We argue by contradiction and assume that $\operatorname{cd}(\Gamma)>4$. Let $\mathcal{A}$ denote the algebraic closure of $p(\Gamma)$. Then $\mathcal{A}$ is a solvable, nonabelian subgroup of ${\operatorname{GL}}_{2}^{+}(\mathbb{R})$ with finitely many connected components. In particular, up to conjugacy, the identity component of $\mathcal{A}$ is contained in the subgroup of upper triangular matrices in ${\operatorname{GL}}_{2}^{+}(\mathbb{R})$, which can be written as $\mathbb{R}\times B$, where
$B=\left\{\begin{pmatrix}e^{t}&x\\ 0&e^{-t}\end{pmatrix}\ \Big|\ t,x\in\mathbb{R}\right\},$
and $\mathbb{R}$ is identified with the subgroup of homotheties. It follows that, up to finite index, the group $p(\Gamma)$ is contained in $\mathbb{R}\times B$. The group $\mathbb{R}\times B$ can be written as $A\ltimes(\mathbb{R}\times U_{1}),$ where $U_{1}$ is the nilradical of $B$ (see (6)) and the $A$-factor corresponds to the hyperbolic one-parameter group given by
$A=\{\mathrm{diag}(e^{t},e^{-t})\mid t\in\mathbb{R}\}.$
Next, we study the possible structure of $p(\Gamma)$.
Lemma 6.9.
If $p(\Gamma)$ is not abelian, then the identity component $\overline{p(\Gamma)}^{\circ}$ coincides with one of the following:
  1. (1)
    the nilradical of $B$, which is $U_{1}$;
  2. (2)
    the nilradical of $\mathbb{R}\times B$, which is $\mathbb{R}\times U_{1}$.
Proof.
We write $\mathbb{R}\times B$ as $A\ltimes(\mathbb{R}\times U_{1})$, and let $\pi\colon A\ltimes(\mathbb{R}\times U_{1})\to A$ be the natural projection, and set $H:=\overline{p(\Gamma)}^{\circ}$. We claim that $\pi(H)$ is necessarily trivial. Suppose, by contradiction, that $\pi(H)$ is nontrivial. Since $\pi(H)$ is a connected subgroup of $A$, it follows that $\pi(H)=A$. We distinguish cases according to the dimension of $H$.
  • If $\dim(H)=1$, then up to conjugacy we may assume that $H$ is a one-parameter subgroup of $A\ltimes\mathbb{R}=A\times\mathbb{R}$. Since $\pi(H)=A$, the group $H$ is not contained in the $\mathbb{R}$-factor, which consists of pure homotheties. It follows that the normalizer of $H$ is abelian. In particular $\overline{p(\Gamma)}$ is abelian, which is a contradiction.
  • If $\dim(H)=2$, then ${\operatorname{Ker}}(\pi)\cap H\subset\mathbb{R}\times U_{1}$ is a one-dimensional subgroup invariant under $\mathrm{Ad}_{h}$ for all $h\in\pi(H)=A$. For $h=\operatorname{diag}(e^{s},e^{-s})\in A$, the adjoint action of $h$ on $\mathbb{R}\times U_{1}$ is given, in the canonical basis of the Lie algebra of $\mathbb{R}\times U_{1}$, by
    $\begin{pmatrix}1&0\\ 0&e^{2s}\end{pmatrix}$. (17)
    Observe that $\pi(H)$ preserves ${\operatorname{Ker}}(\pi)\cap H$, and since $H$ is nilpotent, the action of $\pi(H)$ on ${\operatorname{Ker}}(\pi)\cap H$ is unipotent, equation (17) implies that necessarily ${\operatorname{Ker}}(\pi)\cap H=\mathbb{R}$. Thus we obtain a short exact sequence $1\to\mathbb{R}\to H\to A\to 1$, and hence $H\cong A\ltimes\mathbb{R}=A\times\mathbb{R}$. The normalizer of such a group is $A\times\mathbb{R}$, and therefore $\overline{p(\Gamma)}$ is abelian, again a contradiction.
We conclude that $\pi(H)$ must be trivial and so $H\leq\mathbb{R}\times U_{1}$. We now argue according to the dimension of $H$.
  • If $\dim(H)=1$, then $H$ is a one-parameter subgroup of $\mathbb{R}\times U_{1}$. Since the projection of $H$ onto $A$ is trivial, it follows that either $H=U_{1}$, or the normalizer of $H$ is abelian. Hence $H=U_{1}$.
  • If $\dim(H)=2$, then necessarily $H=\mathbb{R}\times U_{1}$, since $H\leq\mathbb{R}\times U_{1}$.
Next, we show the following.
Lemma 6.10.
If $\operatorname{cd}(\Gamma)>4$, then $\overline{p(\Gamma)}^{\circ}$ cannot be equal to $\mathbb{R}\times U_{1}$.
Proof.
Assume by contradiction that $\overline{p(\Gamma)}^{\circ}=\mathbb{R}\times U_{1}$. Consider the natural projection $\mathbb{R}\times B\to A$. Its kernel is $\mathbb{R}\times U_{1}$, which coincides with $\overline{p(\Gamma)}^{\circ}$. Hence, the projection of $\overline{p(\Gamma)}$ is discrete in $A\cong\mathbb{R}$. If this projection is trivial, then $\overline{p(\Gamma)}$ is abelian, as it coincides with $\mathbb{R}\times U_{1}$. This contradicts the hypothesis of Proposition 6.8. Therefore, $\overline{p(\Gamma)}\cong\mathbb{Z}\ltimes(\mathbb{R}\times U_{1})$.
Let $\Gamma_{2}=\Gamma\cap\big((\mathbb{R}\times U_{1})\ltimes N\big)$, by Theorem C, the group $\Gamma_{2}$ admits a nilpotent syndetic hull $S_{2}$ inside $(\mathbb{R}\times U_{1})\ltimes N$. The projection to the $A$ factor provides the short exact sequence
$1\to\Gamma_{2}\to\Gamma\to\mathbb{Z}\to 1.$
Using the assumption $\operatorname{cd}(\Gamma)>4$, we deduce that $\operatorname{cd}(\Gamma_{2})>3$, and hence $\dim(S_{2})\geq\operatorname{cd}(\Gamma_{2})>3.$ In particular $S_{2}\cap\mathbb{R}^{4}$ is nontrivial and invariant under the projection of $S_{2}$ to the linear part in $\mathbb{R}\ltimes(U_{1}\times U)$. However, the $\mathbb{R}$-action on $S_{2}\cap\mathbb{R}^{4}$ is not unipotent (see (16)), which contradicts the nilpotency of $S_{2}$. This completes the proof. ∎
We now investigate the case where $\overline{p(\Gamma)}^{\circ}$ is the nilradical of $B$.
Proposition 6.11.
If $\operatorname{cd}(\Gamma)>4$, then $\overline{p(\Gamma)}^{\circ}$ cannot be equal to $U_{1}$.
To prove this, we consider the projection $j\colon\mathbb{R}\times B\longrightarrow\mathbb{R}\times A$ modulo $U_{1}$. Since $\overline{p(\Gamma)}^{\circ}={\operatorname{Ker}}(j)$, it follows that $j(\overline{p(\Gamma)})$ is discrete. We define
$\Gamma^{\prime}:=j\bigl(p(\Gamma)\bigr),$ (18)
which is a discrete subgroup of $\mathbb{R}^{2}$. Since $U_{1}\ltimes N$ is an algebraic nilpotent group, the subgroup $\Gamma_{1}:=\Gamma\cap(U_{1}\ltimes N)$ admits a nilpotent syndetic hull, which we denote by $S_{1}$; this is the Malcev closure of $\Gamma_{1}$.
Lemma 6.12.
The Malcev closure $S_{1}\leq U_{1}\ltimes N$ of $\Gamma_{1}:=\Gamma\cap(U_{1}\ltimes N)$ is abelian. Moreover, its projection onto $U_{1}$ is nontrivial.
Proof.
Since $\Gamma_{1},\Gamma_{1}]\leq\Gamma\cap N=\{0\},$ it follows that $\Gamma_{1}$ is abelian. Moreover $\Gamma_{1}$ is a uniform lattice in the simply connected nilpotent Lie group $S_{1}$, which implies that $S_{1}$ itself is abelian.
To prove that the projection of $S_{1}$ onto $U_{1}$ is nontrivial, assume by contradiction that $S_{1}\leq N$. Then $\Gamma_{1}\leq N$, and hence $\Gamma_{1}\leq\Gamma\cap N=\{0\},$ so $\Gamma_{1}$ is trivial. Consequently $\Gamma$ is a subgroup of $\mathbb{R}\times A$, which is abelian, and therefore $p(\Gamma)$ is abelian. This contradicts our standing assumption that $p(\Gamma)$ is not abelian. ∎
We now begin the investigation of $\Gamma^{\prime}$ depending on its rank.
Lemma 6.13.
The subgroup $\Gamma^{\prime}$ in (18) cannot have rank $1$.
Proof.
Assume by contradiction that $\Gamma^{\prime}$ has rank $1$, and set $\Gamma_{1}=\Gamma\cap(U_{1}\ltimes N)$ with syndetic hull $S_{1}$. Using the short exact sequence $1\to\Gamma_{1}\to\Gamma\to\Gamma^{\prime}\to 1$ and the assumption that $\operatorname{cd}(\Gamma)>4$, we deduce that $\operatorname{cd}(\Gamma_{1})>3$ and, in particular $\dim(S_{1})\geq 4$. It follows from Lemma 8.1 that $S_{1}$ is the group of pure translations. This implies that $\Gamma_{1}\cap N$ is nontrivial, which contradicts the fact that $\Gamma\cap N$ is trivial. ∎
We now deal with the case where $\Gamma^{\prime}$ has rank $2$. We state the following result.
Proposition 6.14.
The subgroup $\Gamma^{\prime}$ in (18) cannot have rank $2$.
We begin with the following result.
Lemma 6.15.
If $\Gamma^{\prime}$ has rank $2$, then $\dim(S_{1})=3$, and the group $\Gamma$ admits a syndetic hull $S_{\Gamma}\leq(\mathbb{R}\times B)\ltimes N$ of dimension $5$ such that $S_{\Gamma}\cap(U_{1}\ltimes N)=S_{1}$, where $S_{1}\leq U_{1}\ltimes N$ is the Malcev closure of $\Gamma_{1}=\Gamma\cap(U_{1}\ltimes N)$.
Proof.
Using the short exact sequence $1\to\Gamma_{1}\to\Gamma\to\Gamma^{\prime}\to 1$, we deduce that $5\leq\operatorname{cd}(\Gamma)\leq\operatorname{cd}(\Gamma_{1})+2$, and hence $\operatorname{cd}(\Gamma_{1})\geq 3$. If $\operatorname{cd}(\Gamma_{1})>3$, then $\dim(S_{1})\geq 4$, and, as in the proof of Lemma 6.13, we deduce that $S_{1}$ is the group of pure translations. This implies that $\Gamma\cap N$ is nontrivial, which is a contradiction. Therefore $\operatorname{cd}(\Gamma_{1})=3$, and consequently $\operatorname{cd}(\Gamma)=5$.
Since $\Gamma$ is solvable and $\mathbb{R}\times B\ltimes N$ is an algebraic group, $\Gamma$ admits a syndetic hull $S_{\Gamma}$ by Theorem 2.7. Moreover, since $\mathbb{R}\times B\ltimes N$ is contractible, the group $S_{\Gamma}$ is contractible, and hence $\operatorname{cd}(\Gamma)=\dim(S_{\Gamma})$. Let $j_{1}:(\mathbb{R}\times A)\ltimes(U_{1}\ltimes N)\to\mathbb{R}\times A$ be the natural projection. Note that the restriction $j_{1}|_{S_{\Gamma}}:S_{\Gamma}\to\mathbb{R}\times A$ is surjective. Indeed, $j_{1}(S_{\Gamma})$ is a connected Lie subgroup of $\mathbb{R}\times A\cong\mathbb{R}^{2}$ containing the discrete group $\Gamma^{\prime}$ of rank $2$, and so $j_{1}(S_{\Gamma})=\mathbb{R}\times A$. Since ${\operatorname{Ker}}(j_{1}|_{S_{\Gamma}})=S_{\Gamma}\cap(U_{1}\ltimes N)$, we deduce that $\dim(S_{\Gamma}\cap(U_{1}\ltimes N))=3$.
The group $S_{\Gamma}\cap(U_{1}\ltimes N)$ contains $\Gamma_{1}$, and since $\operatorname{cd}(\Gamma_{1})=\dim(S_{\Gamma}\cap(U_{1}\ltimes N))$, we conclude that $\Gamma_{1}$ is cocompact in $S_{\Gamma}\cap(U_{1}\ltimes N)$. As the Malcev closure of $\Gamma_{1}$ in $U_{1}\ltimes N$ is unique, the result follows. ∎
We arrive at the following result.
Lemma 6.16.
Let $S_{\Gamma}$ and $S_{1}$ as in Lemma 6.15, and let $S_{0}=S_{1}\cap N$ with Lie algebra $\mathfrak{s}_{0}$. Take $g=(\delta,n)\in S_{\Gamma}$, where $\delta=\mathrm{diag}(\lambda\mu,\lambda\mu^{-1})\in\mathbb{R}\times A$. Then ${\operatorname{det}}\bigl(\mathrm{Ad}_{g\mid\mathfrak{s}_{0}}\bigr)=1/\mu^{2}$.
First we record the following basic computation.
Sublemma 6.16.1.
Let $\lambda\in\mathbb{R}$ and $\mu\in\mathbb{R}^{*}$, and consider the diagonal matrix $\delta=\mathrm{diag}(\lambda\mu,\lambda\mu^{-1})$. Then the adjoint action of $\delta$ on $U_{1}\ltimes N$ is given, in the basis $u_{1},u,T_{1},T_{2},T_{3},T_{4}$ (see Table 1), by
$\mathrm{Ad}_{\delta}=\mathrm{diag}\bigl(\mu^{2},\,\lambda^{2},\,\lambda\mu,\,\lambda\mu^{-1},\,\lambda^{-1}\mu^{-1},\,\lambda^{-1}\mu\bigr).$
Proof of Lemma 6.16.
Let $g=(\delta,n)\in S_{\Gamma}$, where $\delta\in\mathbb{R}\times A$ and $n\in N_{1}$. Let $\mathfrak{s}_{1}$ be the Lie algebra of $S_{1}$. We claim that
${\operatorname{det}}\bigl(\mathrm{Ad}_{g\mid\mathfrak{s}_{1}}\bigr)=1.$
Indeed, by Lemma 6.15, we have $S_{1}=S_{\Gamma}\cap(U_{1}\ltimes N)$, and hence $S_{1}$ is normal in $S_{\Gamma}$. This follows from the fact that $U_{1}\ltimes N$ is normal in $(\mathbb{R}\times A)\ltimes(U_{1}\ltimes N)$. In particular, we have the short exact sequence
$1\longrightarrow S_{1}\longrightarrow S_{\Gamma}\longrightarrow\mathbb{R}\times A\longrightarrow 1.$
Since $\Gamma$ is a uniform lattice in $S_{\Gamma}$, then $S_{\Gamma}$ is unimodular, and so we obtain
${\operatorname{det}}\bigl(\mathrm{Ad}_{g\mid\mathfrak{s}_{\Gamma}}\bigr)={\operatorname{det}}\bigl(\mathrm{Ad}_{g\mid\mathfrak{s}_{1}}\bigr){\operatorname{det}}\bigl(\mathrm{Ad}_{g\mid(\mathfrak{s}_{\Gamma}/\mathfrak{s}_{1})}\bigr)=1.$
As $S_{\Gamma}/S_{1}\cong\mathbb{R}\times A$, a direct computation shows that ${\operatorname{det}}\bigl(\mathrm{Ad}_{g\mid(\mathfrak{s}_{\Gamma}/\mathfrak{s}_{1})}\bigr)=1$, and hence ${\operatorname{det}}\bigl(\mathrm{Ad}_{g\mid\mathfrak{s}_{1}}\bigr)=1$.
Next, since $N$ is normal in $(\mathbb{R}\times B)\ltimes N$, we have the short exact sequence
$1\longrightarrow S_{0}\longrightarrow S_{1}\longrightarrow U_{1}\longrightarrow 1.$
Therefore,
${\operatorname{det}}\bigl(\mathrm{Ad}_{g\mid\mathfrak{s}_{1}}\bigr)={\operatorname{det}}\bigl(\mathrm{Ad}_{g\mid\mathfrak{s}_{0}}\bigr){\operatorname{det}}\bigl(\mathrm{Ad}_{g\mid(\mathfrak{s}_{1}/\mathfrak{s}_{0})}\bigr)=1.$
By Lemma 6.16.1, ${\operatorname{det}}\bigl(\mathrm{Ad}_{g\mid(\mathfrak{s}_{1}/\mathfrak{s}_{0})}\bigr)=\mu^{2}$, and the result follows. ∎
Finally, we prove Proposition 6.14.
Proof of Proposition 6.14.
Let $\delta_{1}=\mathrm{diag}(\lambda_{1}\mu_{1},\lambda_{1}\mu_{1}^{-1})$ and $\delta_{2}=\mathrm{diag}(\lambda_{2}\mu_{2},\lambda_{2}\mu_{2}^{-1})$ be the two generators of the group $\Gamma^{\prime}$ defined in (18). Note that
$(\log|\lambda_{1}\mu_{1}|,\log|\lambda_{1}\mu_{1}^{-1}|)\quad\text{and}\quad(\log|\lambda_{2}\mu_{2}|,\log|\lambda_{2}\mu_{2}^{-1}|)$
are linearly independent in $\mathbb{R}^{2}$. In the rest of the proof, we will simply say that $\delta_{1}$ and $\delta_{2}$ are linearly independent.
By Lemma 6.12, the group $S_{1}\leq U_{1}\ltimes N$ is an abelian group of dimension $3$ with a nontrivial projection onto $U_{1}$. It then follows from Lemma 8.2 that the Lie algebra $\mathfrak{s}_{1}$ is one of $\mathfrak{a}_{1}$, $\mathfrak{a}_{3}^{\pm}$, or $\mathfrak{a}_{4}$. Denoting by $\mathfrak{s}_{0}$ the Lie algebra of $S_{1}\cap N$, the proof proceeds according to the different possibilities for $\mathfrak{s}_{1}$.
Case $\mathfrak{s}_{1}=\mathfrak{a}_{1}$. In this case $\mathfrak{s}_{0}={\operatorname{Span}}\{T_{1},T_{4}\}$. Since $T_{1}$ is central in $N$, for $g_{i}=(\delta_{i},n_{i})$ with $\delta_{i}\in\Gamma^{\prime}$ and $n_{i}\in N_{1}$, $i=1,2$, we have $\mathrm{Ad}_{g_{i}}(T_{1})=\mathrm{Ad}_{\delta_{i}}(T_{1})=\lambda_{i}\mu_{i}T_{1}$.
Therefore, by Sublemma 6.16.1, the restriction of $\mathrm{Ad}_{g_{i}}$ to $\mathfrak{s}_{0}$ has the form, in the basis $T_{1},T_{4}$,
$\begin{pmatrix}\lambda_{i}\mu_{i}&\star\\ 0&\lambda_{i}^{-1}\mu_{i}\end{pmatrix}$.
Thus ${\operatorname{det}}(\mathrm{Ad}_{g_{i}\mid\mathfrak{s}_{0}})=\mu_{i}^{2}$, and using Lemma 6.16 we obtain $\mu_{i}^{4}=1$, hence $\mu_{i}=\pm 1$. In particular, $\delta_{1}$ and $\delta_{2}$ are not linearly independent, a contradiction.
Case $\mathfrak{s}_{1}=\mathfrak{a}_{3}^{\pm}$. Here $\mathfrak{s}_{0}={\operatorname{Span}}\{T_{1},T_{2}\pm T_{4}\}$. One checks that
$\mathrm{Ad}_{g_{i}}(T_{2})=\lambda_{i}\mu_{i}^{-1}T_{2}+x_{i}T_{1},\qquad\mathrm{Ad}_{g_{i}}(T_{4})=\lambda_{i}^{-1}\mu_{i}T_{4}+y_{i}T_{1},$ (19)
for some $x_{i},y_{i}\in\mathbb{R}$. For $\mathfrak{s}_{0}$ to be invariant under $\mathrm{Ad}_{g_{i}}$, we must have $\lambda_{i}\mu_{i}^{-1}=\lambda_{i}^{-1}\mu_{i}$, which implies $\lambda_{i}=\pm\mu_{i}$. Hence $\delta_{1}$ and $\delta_{2}$ are not linearly independent, a contradiction.
Case $\mathfrak{s}_{1}=\mathfrak{a}_{4}$. In this case, $\mathfrak{s}_{0}={\operatorname{Span}}\{u+\beta T_{2}+\gamma T_{4},\;T_{1}\}$. The restriction of $\mathrm{Ad}_{g_{i}}$ to $\mathfrak{s}_{0}$ has the following form in the basis $T_{1},\,u+\beta T_{2}+\gamma T_{4}$:
$\begin{pmatrix}\lambda_{i}\mu_{i}&\star\\ 0&\lambda_{i}^{2}\end{pmatrix}$.
Thus ${\operatorname{det}}(\mathrm{Ad}_{g_{i}\mid\mathfrak{s}_{0}})=\mu_{i}\lambda_{i}^{3}$, and by Lemma 6.16 we obtain $\lambda_{i}^{3}\mu_{i}^{3}=1$. In particular $\lambda_{i}=\mu_{i}^{-1}$, and again $\delta_{1}$ and $\delta_{2}$ are not linearly independent, a contradiction. ∎
Proof of Proposition 6.11.
By contradiction, assume that $\overline{p(\Gamma)}^{\circ}=U_{1}$. Consider the group $\Gamma^{\prime}$ defined in (18), so that $\overline{p(\Gamma)}\cong\Gamma^{\prime}\ltimes U_{1}.$ Lemma 6.13 and Proposition 6.14 imply that $\Gamma^{\prime}$ is trivial. Hence $\overline{p(\Gamma)}=U_{1}$ is abelian, which yields a contradiction. ∎
The proof of Proposition 6.8 follows by combining Lemma 6.9, Lemma 6.10, and Proposition 6.11. This also completes the proof of Proposition 6.6.
Proof of Proposition 6.1.
Let $\Gamma$ and $\Omega$ as in Proposition 6.1. Let $T=q(\Gamma)\cap\left(\{\mathrm{Id}\}\times\mathbb{R}^{2}\right)$ be the group of pure translations in $q(\Gamma)$ as in (14). If $T$ is not discrete, then completeness follows from Corollary 4.2. If $T$ is discrete, then by Lemma 6.2, the group $T$ is either trivial or isomorphic to $\mathbb{Z}$. In both cases, completeness follows from Propositions 6.3 and 6.6. This completes the proof. ∎
We now have all the tools to prove our main theorem.
Proof of Theorem A.
Let $\Gamma$ and $\Omega$ be as in Theorem A. Then, up to passing to a subgroup of index at most two, we may assume that $\Gamma\leq{\operatorname{SO}}_{0}(2,2)\ltimes\mathbb{R}^{2,2}$. We then argue according to whether $q:\Gamma\to{\operatorname{GL}}_{2}^{+}(\mathbb{R})\ltimes_{\mathsf{t}}\mathbb{R}^{2}$ is injective or not. In both cases, we have completeness from Propositions 5.1 and 6.1. ∎
7. Proof of Theorem C
The goal of this section is to prove Theorem C. Let $G$ be a homothety lie group and $G_{\theta}=R\ltimes_{\theta}G$ as in Theorem C. The first step is to show that $G_{\theta}$ is linear.
Lemma 7.1.
The group $G_{\theta}=R\ltimes_{\theta}G$ is linear, i.e., there exists an injective homomorphism $G_{\theta}\hookrightarrow{\operatorname{GL}}_{n}(\mathbb{C})$ for some $n\in\mathbb{N}$.
Proof.
The idea of the proof is inspired from [HKMZ25, Appendix B]. Consider the natural morphism $f:G_{\theta}\to{\sf{Aut}}(G)\ltimes G,\quad f(r,g)=(\theta(r),g).$ Note that this morphism is not necessarily injective. We claim that the group ${\sf{Aut}}(G)\ltimes G$ is linear. Indeed, since $G$ admits homotheties, the center of ${\sf{Aut}}(G)\ltimes G$ is trivial. Hence the adjoint representation $\mathrm{Ad}:{\sf{Aut}}(G)\ltimes G\to{\operatorname{GL}}(\mathfrak{l})$ is faithful, where $\mathfrak{l}$} is the Lie algebra of ${\sf{Aut}}(G)\ltimes G$. Now define $\Phi:G_{\theta}\to R\times{\operatorname{GL}}({\mathfrak{l}}),\ \Phi(r,g)=(r,\mathrm{Ad}(f(r,g))).$ Then $\Phi$ is clearly a faithful morphism into $R\times{\operatorname{GL}}({\mathfrak{l}})$, which is a linear group. The claim follows. ∎
We will now recall the following well known result.
Lemma 7.2 (Strong Zassenhaus Lemma, Theorem 4.1.7 [Thu14], Proposition 8.16 [Rag72]).
Let $G$ be a Lie group. There exists a neighborhood $V$ of $1$ in $G$ such that any discrete subgroup $\Gamma$ of $G$ generated by $V\cap\Gamma$ admits a nilpotent syndetic hull $S$ in $G$.
We recall also the following lemma.
Lemma 7.3.
[HKMZ25, Lemma 4.3] Let $\Lambda$ be a subgroup of a Lie group $G$. Define $\Lambda_{0}:=\Lambda\cap{\overline{\Lambda}}^{\mathsf{o}}$, where ${\overline{\Lambda}}^{\mathsf{o}}$ denotes the identity component of the topological closure of $\Lambda$. Then the subgroup $\Lambda_{0}$ can be generated by $\Lambda_{0}\cap V^{\prime}$ for any neighborhood $V^{\prime}$ of identity in ${\overline{\Lambda}}^{\mathsf{o}}$.
The last result needed to prove Theorem C, is the following.
Lemma 7.4.
[CD89, Lemma 1.3.2] A discrete subgroup of ${\operatorname{GL}}_{n}(\mathbb{C})$ which is locally nilpotent (i.e. any finitely generated subgroup is nilpotent) is nilpotent.
Proof of Theorem C.
As in the statement let $\Gamma$ be a discrete subgroup of $G_{\theta}$. Let $H$ be the closure of $\pi(\Gamma)$. We consider the non-discrete part of $\Gamma$, namely $\Gamma_{nd}=\Gamma\cap\pi^{-1}(H^{\circ})$. By Lemma 7.1, the group $\Gamma_{nd}$ is linear. The goal is to show that every finitely generated subgroup $\Gamma_{0}$ of $\Gamma_{nd}$ is nilpotent and then conclude using Lemma 7.4 that $\Gamma_{nd}$ is nilpotent. We choose $V_{1}\subset R$ and $V_{2}\subset G$ such that $V_{1}\times V_{2}$ is a strong Zassenhaus neighborhood in $G_{\theta}=R\ltimes_{\theta}G$. Lemma 7.3 applied to $\Lambda=\pi(\Gamma_{0})$ yields that $\Lambda_{0}=\Lambda\cap\overline{\Lambda}^{\circ}$ is generated by $\Lambda_{0}\cap V_{1}$. Let $rh$ be one of finitely many generators of $\Gamma_{0}$, where $r\in R$, and $h\in G$. Then $r=\lambda_{1}\cdot\ldots\cdot\lambda_{k}$ for $\lambda_{j}\in\Lambda_{0}\cap V_{1}$, $j=1,\dots,k$. Choose elements $\gamma_{j}\in\Gamma_{0}$ such that $r(\gamma_{j})=\lambda_{j}$, $j=1,\dots,k$. Then $(\gamma_{1}\cdots\gamma_{k})^{-1}(rh)\in{\operatorname{Ker}}(\pi)\cap\Gamma_{0}=G\cap\Gamma_{0}\penalty 10000\$and so
$rh=\lambda_{1}\cdot\ldots\cdot\lambda_{k}h=\gamma_{1}\cdot\ldots\cdot\gamma_{k}h^{\prime}$
for some $h^{\prime}\in\Gamma_{0}\cap G$. Thus we may replace $rh$ by $\gamma_{1},\dots,\gamma_{k}$ and $h^{\prime}$. Repeating this procedure for every generator of $\Gamma_{0}$, we obtain a set of generators $\{r_{i}h_{i}\}_{i=1}^{m}$, where $r_{i}\in V_{1}$ and $h_{i}\in G$. Let $\Psi$ be a homothety of $G$ that commutes with $\theta$. Then $\Psi$ extends to an automorphism of $G_{\theta}$ by $\Psi_{\theta}=(\mathrm{Id},\Psi),$ whose restriction to $R$ is the identity. Consider $n$ big enough so that $\Psi(h_{i})\in V_{2}$ for all $i=1,\dots,m$. Then $\Psi_{\theta}^{n}(\Gamma_{0})=\langle r_{i}\Psi^{n}(h_{i})\mid\ i=1,\dots,m\rangle$ is generated by elements of $V_{1}\times V_{2}$, thus it is nilpotent by Lemma 7.2. Consequently, also $\Gamma_{0}$ is nilpotent. As $\Gamma_{nd}$ is linear, we conclude that $\Gamma_{nd}$ itself is nilpotent.
Since $\Gamma_{nd}$ is discrete and nilpotent it is in particular, polycyclic [Rag72, Proposition 3.8] so it is finitely generated. Hence, the argument above applies to $\Gamma_{nd}$ itself. Then $\Psi_{\theta}^{n}(\Gamma_{nd})$ for some big $n$ is generated by elements that belongs to the strong Zassenhaus neighborhood of $V_{1}\times V_{2}$. Thus applying Lemma 7.2 gives rise a syndetic hull for $\Gamma_{nd}$.
The last part of the Theorem is to show that $\overline{\pi(\Gamma_{nd})}=H^{\circ}=\pi(S)$, where $S$ is a syndetic hull of $\Gamma_{nd}$ in $H^{\circ}\ltimes G$. By definition $H^{\circ}=\overline{\pi(\Gamma_{nd})}$, on the other hand, $\pi(\Gamma_{nd})\subset\pi(S)$. We have that $\pi(S)\subset H^{\circ}$. Therefore, $H^{\circ}=\overline{\pi(\Gamma_{nd})}\subset\overline{\pi(S)}\subset H^{\circ}$. The claim follows. ∎
8. Appendix: Abelian Subgroups
In this appendix, we record some results concerning the structure of abelian subgroups of $U_{1}\ltimes N$. These results are used throughout the paper in Sections 5 and 6. We rewrite $U_{1}\ltimes N$ as $(U_{1}\times U)\ltimes\mathbb{R}^{4}$. Recall from the description of $U_{1}\times U$ in terms of $(4,4)$ matrices that the Lie algebra $\mathfrak{u}_{1}\times\mathfrak{u}$ is given by
$\left\{\alpha_{x,y}:=\begin{pmatrix}0&y&0&x\\ 0&0&-x&0\\ 0&0&0&0\\ 0&0&-y&0\end{pmatrix}\ \Bigg|\ x,y\in\mathbb{R}\right\},$
so that $\alpha_{1,0}=u$ and $\alpha_{0,1}=u_{1}$.
Let $G$ be an abelian subgroup of $U_{1}\ltimes N$. Let $\mathfrak{g}$} and ${\mathfrak{n}}_{1}$ denote the Lie algebras of $G$ and $U_{1}\ltimes N$, respectively, and let $\ell:{\mathfrak{n}}_{1}\to\mathfrak{u}_{1}\times\mathfrak{u}$ be the projection at the level of Lie algebras of $U_{1}\ltimes N$. In the case where $\ell({\mathfrak{g}})$ is nonzero, the adjoint action $\mathrm{ad}_{\ell({\mathfrak{g}})}$ acts trivially on ${\mathfrak{g}}_{0}:={\mathfrak{g}}\cap\mathbb{R}^{4}$, that is,
${\mathfrak{g}}_{0}\subset\bigcap_{\alpha\in\ell({\mathfrak{g}})}{\operatorname{Ker}}(\mathrm{ad}_{\alpha}|_{\mathbb{R}^{4}}),$
which follows from the fact that $\mathfrak{g}$} is abelian. Next, observe that the restriction of $\mathrm{ad}_{\alpha_{x,y}}$ to $\mathbb{R}^{4}$ coincides with $\alpha_{x,y}$, and hence
${\mathfrak{g}}_{0}\subset\bigcap_{\alpha_{x,y}\in\ell({\mathfrak{g}})}{\operatorname{Ker}}(\alpha_{x,y}).$ (20)
We start with the following result.
Lemma 8.1.
Let $G$ be an abelian subgroup of $U_{1}\ltimes N$ of dimension greater than $4$. Then $G$ is the subgroup of pure translations.
Proof.
Let $G$ be an abelian subgroup of $U_{1}\ltimes N$, and assume by contradiction that it is not contained in the subgroup of pure translations. By assumption, $\ell({\mathfrak{g}})$ is nonzero. Since $\ell({\mathfrak{g}})\cong{\mathfrak{g}}/{\mathfrak{g}}_{0}$, if we assume that $\dim\ell({\mathfrak{g}})=1$, then $\dim({\mathfrak{g}}_{0})\geq 3$. However, if we denote by $\alpha_{x,y}$ a generator of $\ell({\mathfrak{g}})$, then by (20) we have ${\mathfrak{g}}_{0}\subset{\operatorname{Ker}}(\alpha_{x,y})$. Since $x,y\neq 0$, we have $\dim({\operatorname{Ker}}(\alpha_{x,y}))\leq 2$, and hence $\dim({\mathfrak{g}}_{0})\leq\dim({\operatorname{Ker}}(\alpha_{x,y}))\leq 2$, which is a contradiction.
Therefore, we must have $\dim\ell({\mathfrak{g}})=2$, and hence $\dim({\mathfrak{g}}_{0})\geq 2$. Let $\alpha_{0,1}$ and $\alpha_{1,0}$ be generators of $\ell({\mathfrak{g}})$. Then, by (20), we obtain ${\mathfrak{g}}_{0}\subset{\operatorname{Ker}}(\alpha_{0,1})\cap{\operatorname{Ker}}(\alpha_{1,0})={\operatorname{Span}}\{T_{1}\}.$ This contradicts the fact that $\dim{\mathfrak{g}}_{0}\geq 2$, and the proof is complete. ∎
The next result deals with three–dimensional subgroups.
Lemma 8.2.
Let $G$ be an abelian Lie subgroup of $U_{1}\ltimes N$ of dimension $3$ which is not contained in the group of pure translations. Let $u,u_{1},T_{1},T_{2},T_{3},T_{4}$ be the basis of the Lie algebra as in (1). Then the Lie algebra of $G$ is given by
$(1)$ ${\mathfrak{a}}_{1}={\operatorname{Span}}(u_{1}+t,\;T_{1},\;T_{4}),\qquad$ $t\in{\operatorname{Span}}(T_{2},T_{3}),$
$(2)$ ${\mathfrak{a}}_{2}={\operatorname{Span}}(u+t,\;T_{1},\;T_{2}),\qquad$ $t\in{\operatorname{Span}}(T_{3},T_{4}),$
$(3)$ ${\mathfrak{a}}_{3}^{\pm}={\operatorname{Span}}(u\mp u_{1}+t,\;T_{1},\;T_{2}\pm T_{4}),\qquad$ $t\in{\operatorname{Span}}(T_{3},T_{4}),$
$(4)$ ${\mathfrak{a}}_{4}(\alpha,\beta,\gamma)={\operatorname{Span}}(u_{1}+\alpha T_{2}+\beta T_{4},\;u+\beta T_{2}+\gamma T_{4},\;T_{1}),\qquad$ $\alpha,\beta,\gamma\in\mathbb{R},$
Proof.
We proceed with the proof according to the dimension of $\ell({\mathfrak{g}})$, which is nonzero by hypothesis.
Case 1: $\dim(\ell({\mathfrak{g}}))=1$. Then $\ell({\mathfrak{g}})={\operatorname{Span}}(\alpha_{x,y})$ for some $x,y\in\mathbb{R}$. In particular $\dim({\mathfrak{g}}_{0})=2$, hence ${\mathfrak{g}}_{0}={\operatorname{Ker}}(\alpha_{x,y})={\operatorname{Span}}(T_{1},\,xT_{2}-yT_{4})$. Thus
${\mathfrak{g}}={\operatorname{Span}}(X+t,\,T_{1},\,xT_{2}-yT_{4}),\quad X\in\ell({\mathfrak{g}}),\ t\in\mathbb{R}^{4}.$
We observe that $x\neq\pm y$ then ${\operatorname{Ker}}\alpha_{x,y}$ is exactly $T_{1}$, so we exclude this case, since $\dim{\mathfrak{g}}_{0}=2$. In the case $x=y$, then ${\operatorname{Ker}}\alpha_{x,x}={\operatorname{Span}}\{T_{1},T_{2}-T_{4}\}$ So, in this case we have ${\mathfrak{g}}:={\operatorname{Span}}\{u_{1}+u+t,T_{1},T_{2}-T_{4}\}$, moreover, we may assume that $t\in{\operatorname{Span}}\{T_{3},T_{4}\}$. Indeed $t=aT_{2}+bT_{3}+cT_{4}$, then $u_{1}-u+t-a(T_{2}+T_{4})=u_{1}-u+bT_{3}+(c-a)T_{4}=u_{1}-u+t^{\prime}$. In case x=yx=-y, then ${\operatorname{Ker}}\alpha_{x,-x}={\operatorname{Span}}(T_{1},T_{2}+T_{4})$ and so in this case ${\mathfrak{g}}={\operatorname{Span}}\{u_{1}-u+t,T_{1},T_{2}+T_{4}\}$ for $t\in{\operatorname{Span}}\{T_{3},T_{4}\}$.
Case 2: $\dim(\ell({\mathfrak{g}}))=2$. Then $\ell({\mathfrak{g}})=\mathfrak{u}_{1}\times\mathfrak{u}$ and $\dim({\mathfrak{g}}_{0})=1$. By the condition (20),
${\mathfrak{g}}_{0}\subset{\operatorname{Ker}}(\alpha_{1,0})\cap{\operatorname{Ker}}(\alpha_{0,1})={\operatorname{Span}}(T_{1}),$
so ${\mathfrak{g}}_{0}={\operatorname{Span}}(T_{1})$. Hence ${\mathfrak{g}}={\operatorname{Span}}(u+v,\,u_{1}+w,\,T_{1})$ for $v,w\in{\operatorname{Span}}(T_{2},T_{3},T_{4})$. Write
$v=pT_{2}+qT_{3}+rT_{4},\qquad w=aT_{2}+bT_{3}+cT_{4}.$
Since $\mathfrak{g}$} is abelian, we have $u+v,\;u_{1}+w]=0$, which implies $b=q=0\penalty 10000\$and cp=0.c-p=0. Therefore
$v=cT_{2}+rT_{4},\qquad w=aT_{2}+cT_{4},$
for some $a,c,r\in\mathbb{R}$, and ${\mathfrak{g}}={\operatorname{Span}}\big(u_{1}+aT_{2}+cT_{4},u+cT_{2}+rT_{4},\;T_{1}\big)$. Hence, for $a=\alpha$, $c=\beta$ and $r=\gamma$, the claim follows. ∎
The last result needed concerns abelian subgroups of $N$ of dimension $2$.
Lemma 8.3.
Let $G$ be an abelian subgroup of dimension two of $N=U\ltimes\mathbb{R}^{4}$ which is not contained in the group of pure translations. Then the Lie algebra of $G$ is given by
${\mathfrak{g}}={\operatorname{Span}}(u+t,\;t^{\prime}),$
for $t\in{\operatorname{Span}}(T_{3},T_{4})$ and $t^{\prime}\in{\operatorname{Span}}(T_{1},T_{2})$.
Proof.
By assumption, $\ell({\mathfrak{g}})$ is nonzero and hence $\ell({\mathfrak{g}})={\operatorname{Span}}(\alpha_{1,0})$. In particular, ${\operatorname{Ker}}(\alpha_{1,0})={\operatorname{Span}}(T_{1},T_{2})$, and therefore ${\mathfrak{g}}_{0}\subset{\operatorname{Span}}(T_{1},T_{2})$. It follows that
${\mathfrak{g}}={\operatorname{Span}}(u+t,\;t^{\prime})\qquad t\in\mathbb{R}^{4},\;t^{\prime}\in{\operatorname{Span}}(T_{1},T_{2}),$
and so we may assume that $t\in{\operatorname{Span}}(T_{3},T_{4})$. This completes the proof. ∎

### v2

@@
 This fact is established through the more general Proposition B below.
-Let $G\leq{\operatorname{SL}}(d,\mathbb{R})$ be a semisimple Lie subgroup, and fix a Cartan decomposition $G=KA^{+}K$, where $K\leq G$ is a maximal compact subgroup and $A^{+}$ is the exponential of a closed positive Weyl chamber (see Section 3.1).
-For a subgroup $G^{\prime}\leq G$, denote by $A^{+}(G^{\prime})$ the set of Cartan projections of elements of $G^{\prime}$ with respect to this decomposition.
-Fix once and for all a Euclidean norm $|\cdot|$ on $\mathbb{R}^{d}$ that is invariant under the maximal compact subgroup $K$.
-We say that $A^{+}(G^{\prime})$ does not contract a subspace $F\subset\mathbb{R}^{d}$ with respect to $|\cdot|$ if $|g\cdot x|\geq|x|$ for all $g\in A^{+}(G^{\prime})$ and all $x\in F$.
-Fix $p\in\llbracket 1,d\rrbracket$, and consider $\operatorname{Gr}_{p}(\mathbb{R}^{d})$, the Grassmannian of $p$-dimensional subspaces.
-We say that $E,F\in\operatorname{Gr}_{p}(\mathbb{R}^{d})$ are transverse if $E+F=\mathbb{R}^{d}$ (in particular $p\geq\tfrac{d}{2}$).
-We prove the following result.
+Fix once and for all a Euclidean norm $|\cdot|$ on $\mathbb{R}^{d}$.
+We say that $g\in\mathrm{SL}_{d}(\mathbb{R})$ does not contract a subspace $F\subset\mathbb{R}^{d}$ with respect to $|\cdot|$ if $|g\cdot x|\geq|x|$ for all $x\in F$.
 Proposition B.
-Let $\Omega$ be a domain in $\mathbb{R}^{d}$, different from $\mathbb{R}^{d}$, and let $\Gamma\leq{\operatorname{SL}}_{d}(\mathbb{R})\ltimes\mathbb{R}^{d}$ be a discrete group acting freely, properly discontinuously, and cocompactly on $\Omega$.
-Suppose that the linear part $L(\Gamma)$ of $\Gamma$ is contained in a semisimple Lie group $G\leq{\operatorname{SL}}_{d}(\mathbb{R})$ such that:
+Let $\Omega$ be a domain in $\mathbb{R}^{d}$, different from $\mathbb{R}^{d}$, and let $\mathsf{H}\leq{\operatorname{SL}}_{d}(\mathbb{R})\ltimes\mathbb{R}^{d}$ be a group acting properly and cocompactly on $\Omega$.
+Let us denote by $L(\mathsf{H})$ the linear part of $\mathsf{H}$.
+Assume there exists $1\leq p\leq d-1$ and a closed $L(\mathsf{H})$-invariant subset $\mathcal{F}$ of $\operatorname{Gr}_{p}(\mathbb{R}^{d})$, such that:
   1. (1)
-    there exists $\frac{d}{2}\leq p\leq d$ such that $G$ preserves an orbit $\mathcal{F}\subset\operatorname{Gr}_{p}(\mathbb{R}^{d})$ whose closure consists of pairwise transverse subspaces;
+    for all $g\in L(\mathsf{H})$, there exists an element of $\mathcal{F}$ which is not contracted by $g$;
   2. (2)
-    there exists $H\in\mathcal{F}$ such that $A^{+}(L(\Gamma))$ does not contract $H$.
-Then there exists $H^{\prime}$ in the closure of $\mathcal{F}$ that is $L(\Gamma)$-invariant, and $\Omega$ is foliated by affine subspaces parallel to $H^{\prime}$.
-This proposition generalizes a result of Tholozan [Tho15, Proposition 2.4], where the statement is proved in the case $G=\mathrm{U}(n,1)$ and $\mathcal{F}$ is the orbit of a complex hyperplane.
-In our case, the set of isotropic planes in $\mathbb{R}^{2,2}$ has two connected components, each of which is an orbit under the identity component ${\operatorname{SO}}_{0}(2,2)$; moreover, the elements of each orbit are pairwise transverse (see Proposition 3.7).
+    If $x,y\in\mathcal{F}$ are distinct, then $x+y=\mathbb{R}^{d}$.
+Then there exists $F$ in $\mathcal{F}$ that is $L(\mathsf{H})$-invariant, and $\Omega$ is foliated by affine subspaces parallel to $F$.
+In this proposition, the existence of an element in $\operatorname{Gr}_{p}(\mathbb{R}^{d})$ which is non-contracted by $L(\mathsf{H})$ suggests that $n-p\geq\operatorname{disc}(L(\mathsf{H}))$, where $\operatorname{disc}(L(\mathsf{H}))$ is the disompacity of $L(\mathsf{H})$ in the sense of Carrière [Car89].
+In Carrière’s paper, the assumption of discompacity 1 is necessary.
+Proposition B says that, by adding a transversality assumption, we can relax this condition and get an intermediate structural result for groups of potentially higher dicsompacity.
+It generalizes a result of Tholozan [Tho15, Proposition 2.4], where the statement is proved in the case where $L(\mathsf{H})$ is contained in $\mathrm{U}(n,1)$ and $\mathcal{F}$ is the set of a complex hyperplanes.
+In our case, the set of isotropic planes in $\mathbb{R}^{2,2}$ has two connected components, each of which is an orbit under the identity component ${\operatorname{SO}}_{0}(2,2)$; moreover, the elements of each orbit are pairwise transverse (see Proposition 3.9), meaning that if $x,y$ are two distinct isotropic planes in the orbit, then $x+y=\mathbb{R}^{4}$, i.e.
+they satisfy the second condition of Proposition B.
 Let us briefly explain this well known fact.
@@
 Indeed, their intersection coincides with the set of fixed vectors of $AB^{-1}$, and a nontrivial element of ${\operatorname{SO}}(2)$ has no nonzero fixed vectors.
-This property no longer holds in ${\operatorname{SO}}(n)$ for $n>2$, and it is precisely at this point that the signature $(2,2)$ plays a crucial role.
-Nevertheless, we believe that Proposition B is of independent interest and may apply in other settings, which deserve further investigation.
-For instance, one may consider the case where $G={\operatorname{SO}}(4,\mathbb{C})$, as well as the case where $G$ is a rank-one irreducible simple Lie subgroup of ${\operatorname{SL}}_{d}(\mathbb{R})$.
-These cases are part of an ongoing project of the three authors.
+This property no longer holds in ${\operatorname{SO}}(d)$ for $d>2$, and it is precisely at this point that the signature $(2,2)$ plays a crucial role.
+Nevertheless, Proposition B applies in other settings, including cases that deserve further investigation.
+For instance, when $d=8$ and $L(\mathsf{H})$ is contained in ${\operatorname{SO}}(4,\mathbb{C})$, or for any $d$ when $L(\mathsf{H})$ is contained in an irreducible simple rank-one Lie subgroup of ${\operatorname{SL}}_{d}(\mathbb{R})$.
+A more general Lie-theoretic corollary of Proposition B, containing these two examples, is stated in Corollary 3.7.
+In another direction, an asymptotic version of Proposition B is part of a work in progress and applies to some cases where $L(\mathsf{H})$ is an $\{\alpha_{p}\}$-Anosov subgroup of ${\operatorname{SL}}_{d}(\mathbb{R})$ in the sense of [Lab06, GW12].
 Having established this, Proposition B applies in our setting and yields a foliation of any domain $\Omega\subset\mathbb{R}^{2,2}$ as in Theorem A by parallel isotropic planes.
@@
 Set $\Gamma_{nd}:=\Gamma\cap\pi^{-1}(H^{\circ})$.
-Then $\Gamma_{nd}$ admits a nilpotent syndetic hull $S$ in $G_{\theta}$ with the property that $\overline{\pi(\Gamma_{nd})}=\overline{\pi(N)}=H^{\circ}$.
+Then $\Gamma_{nd}$ admits a nilpotent syndetic hull $S$ in $G_{\theta}$ with the property that $\overline{\pi(\Gamma_{nd})}=\overline{\pi(S)}=H^{\circ}$.
 A syndetic hull of $\Gamma_{nd}$ is a closed connected subgroup containing $\Gamma_{nd}$ as a uniform lattice.
@@
 Theorem 2.3 ([GH84, Cor. 2.11]).
-Let $\Omega\subset\mathbb{R}^{d}$ be an open set divided by a discrete subgroup $\Gamma\leq{\operatorname{SL}}(d,\mathbb{R})\ltimes\mathbb{R}^{d}$. Then $\operatorname{cd}(\Gamma)\geq d$.
+Let $\Omega\subset\mathbb{R}^{d}$ be an open set divided by a discrete subgroup $\Gamma\leq{\operatorname{SL}}_{d}(\mathbb{R})\ltimes\mathbb{R}^{d}$. Then $\operatorname{cd}(\Gamma)\geq d$.
 For instance, this implies that a free group cannot divide an open set in $\mathbb{R}^{d}$ when $d\geq 2$, since the cohomological dimension of such group is $1$.
@@
 Then the linear part of $\Gamma$ preserves a $2$-dimensional totally isotropic plane $P_{0}\subset\mathbb{R}^{2,2}$, and $\Omega$ is foliated by affine translates of $P_{0}$.
-3.1. Cartan projection
-An important feature of the proof of Proposition 3.1 is the Cartan projection.
-Let $G$ be a semisimple Lie group, and denote by $\mathfrak{g}$ its Lie algebra.
-Let $K\leq G$ be a maximal compact subgroup and $\mathfrak{h}$ be the orthogonal of the Lie algebra $\mathfrak{k}$ of $K$ in $\mathfrak{g}$} for the Killing form.
-A Cartan subspace $\mathfrak{a}$ of $\mathfrak{g}$} is a maximal abelian subalgebra of $\mathfrak{h}$.
-The restricted Weyl group $W$ of $G$ is the quotient of the normalizer of $\mathfrak{a}$ in $K$ (for the adjoint action) by the centralizer of $\mathfrak{a}$ in $K$.
-A fundamental system of $\mathfrak{a}$ for the action of $W$ is called a closed Weyl chamber.
-It is a closed convex cone with nonempty interior.
-Given such a closed Weyl chamber $\overline{\mathfrak{a}}^{+}$, we denote by $A^{+}$ the exponential of $\overline{\mathfrak{a}}^{+}$ ; the group $G$ then admits the following Cartan decomposition
-$G=KA^{+}K.$
-Given an element $g\in G$, we can thus write $g=kak^{\prime}$ according to this decomposition.
-The elements $k,k^{\prime}\in K$ are not unique, but the element $a\in A^{+}$ is.
-The Cartan projection of $g$ is the unique element $X\in\overline{\mathfrak{a}}^{+}$ such that $a=\exp(X)$.
 3.1. General reduction result
-For a semisimple Lie subgroup $G\leq{\operatorname{SL}}(d,\mathbb{R})$, we write $G=KA^{+}K$ for a chosen Cartan decomposition of $G$ (in the notation of Section 3.1 above).
-For a subgroup $G^{\prime}\leq G$, we denote by $A^{+}(G^{\prime})$ the set of Cartan projections of elements of $G^{\prime}$ in this decomposition.
-Fix once and for all a Euclidean norm $|\cdot|$ on $\mathbb{R}^{d}$ that is invariant under the maximal compact subgroup $K$.
-We say that a subset $X$ of $G$ does not contract $F$ with respect to $|\cdot|$ if $F$ is invariant under all elements of $X$ and if $|g\cdot x|\geq|x|$ for all $g\in X$ and all $x\in F$.
-In what follows, we will simply say that $X$ does not contract $F$.
+Fix once and for all the Euclidean norm $|\cdot|$ on $\mathbb{R}^{d}$.
+We say that an element $g$ of ${\operatorname{SL}}_{d}(\mathbb{R})$ does not contract $F$ with respect to $|\cdot|$ if $|g\cdot x|\geq|x|$ for all $g\in X$ and all $x\in F$.
+In what follows, we will simply say that $g$ does not contract $F$.
 To prove Proposition 3.1, we first establish the stronger result given by Proposition B, inspired by arguments from [Tho15].
 To fix notation, let $p\in\llbracket 1,d\rrbracket$, and consider $\operatorname{Gr}_{p}(\mathbb{R}^{d})$, the Grassmannian of $p$-dimensional subspaces, note that the linear action of $\mathrm{SL}(d,\mathbb{R})$ induces a natural action on $\operatorname{Gr}_{p}(\mathbb{R}^{d})$.
-We say that $E,F\in\operatorname{Gr}_{p}(\mathbb{R}^{d})$ are transverse if $E+F=\mathbb{R}^{d}$.
 The proof of Proposition B requires some preparation.
 Throughout this section, we work under the assumptions of Proposition B.
-Write $\Omega=\Gamma\cdot\mathcal{C}$ with $\mathcal{C}\subset\Omega$ compact.
+Write $\Omega=\mathsf{H}\cdot\mathcal{C}$ with $\mathcal{C}\subset\Omega$ compact.
 Let $\varepsilon>0$ be such that $\mathcal{C}+B_{\varepsilon}\subset\Omega$, where $B_{\varepsilon}$ denotes the Euclidean ball of radius $\varepsilon$ centered at 0, defined with respect to the Euclidean norm $\lvert\cdot\rvert$.
@@
 Let $y\in\partial\Omega$.
-Then there exist sequences $(g_{n})_{n\in\mathbb{N}}\subset\Gamma$, $(x_{n})_{n\in\mathbb{N}}\subset\mathcal{C}$, and $(H_{n})_{n\in\mathbb{N}}\subset\mathcal{F}$ such that, up to extraction of a subsequence:
+Then there exist sequences $(g_{n})_{n\in\mathbb{N}}\subset\mathsf{H}$, $(x_{n})_{n\in\mathbb{N}}\subset\mathcal{C}$, and $(F_{n})_{n\in\mathbb{N}}\subset\mathcal{F}$ such that, up to extraction of a subsequence:
   1. (1)
@@
   2. (2)
-    for every $n$, one has $(H_{n}+g_{n}\cdot x_{n})\cap B(g_{n}\cdot x,\varepsilon)\subset\Omega$;
+    for every $n$, one has $(F_{n}+g_{n}\cdot x_{n})\cap B(g_{n}\cdot x,\varepsilon)\subset\Omega$;
   3. (3)
-    $H_{n}$ converges in $\operatorname{Gr}_{p}(\mathbb{R}^{n})$ to some element $H_{y}$ in the closure of $\mathcal{F}$, with $(H_{y}+y)\cap B(y,\varepsilon)\subset\partial\Omega$, where $B(y,\varepsilon)$ is Euclidean ball of radius $\varepsilon$ centered at $y$.
+    $F_{n}$ converges in $\operatorname{Gr}_{p}(\mathbb{R}^{d})$ to some element $F^{y}$ in the closure of $\mathcal{F}$, with $(F^{y}+y)\cap B(y,\varepsilon)\subset\partial\Omega$, where $B(y,\varepsilon)$ is Euclidean ball of radius $\varepsilon$ centered at $y$.
 Proof.
-Since $\Omega=\Gamma\cdot\mathcal{C}$, there exist sequences $(x_{n})_{n\in\mathbb{N}}\subset\mathcal{C}$ and $(g_{n})_{n\in\mathbb{N}}\subset\Gamma$ such that $g_{n}\cdot x_{n}\to y$.
+Since $\Omega=\mathsf{H}\cdot\mathcal{C}$, there exist sequences $(x_{n})_{n\in\mathbb{N}}\subset\mathcal{C}$ and $(g_{n})_{n\in\mathbb{N}}\subset\mathsf{H}$ such that $g_{n}\cdot x_{n}\to y$.
 This proves the first point.
@@
 $g_{n}\cdot(x_{n}+B_{\varepsilon})=y_{n}+L(g_{n})\cdot B_{\varepsilon}\subset\Omega.$ (1)
-For each $n$, write the Cartan decomposition $L(g_{n})=w_{n}a_{n}w_{n}^{\prime}$, with $w_{n},w_{n}^{\prime}\in K$ and $a_{n}\in A^{+}(L(\Gamma))$.
-Let $H$ be the $p$-subspace appearing in the hypothesis of Proposition B, and set $H_{n}:=w_{n}\cdot H$.
-We claim that $L(g_{n})\cdot B_{\varepsilon}$ contains $H_{n}\cap B_{\varepsilon}$.
-Indeed, let $v\in w_{n}\cdot H\cap B_{\varepsilon}$ and set $x=w_{n}^{-1}v\in H$.
-Since $K$ preserves the Euclidean norm $\lvert\cdot\rvert$ and $a_{n}$ does not contract $H$, we have $\lvert a_{n}^{-1}x\rvert\leq\lvert x\rvert=\lvert v\rvert\leq\varepsilon$.
-Therefore $u:=w_{n}^{\prime-1}(a_{n}^{-1}x)$ belongs to $B_{\varepsilon}$, and $L(g_{n})u=w_{n}a_{n}w_{n}^{\prime}\bigl(w_{n}^{\prime-1}a_{n}^{-1}x\bigr)=w_{n}x=v$.
-Hence $w_{n}\cdot H\cap B_{\varepsilon}\subset L(g_{n})\cdot B_{\varepsilon}$, which proves the claim.
-It follows from (1) that $\Omega$ contains $y_{n}+(H_{n}\cap B_{\varepsilon})$, establishing the second point.
-By compactness of $\overline{\mathcal{F}}$, we may assume (up to extracting a subsequence) that $H_{n}$ converges in the Grassmannian to some $H_{y}\in\overline{\mathcal{F}}$.
-Then $\overline{\Omega}$ contains $y+(H_{y}\cap B_{\varepsilon})$, and since $y\in\partial\Omega$, the inclusion must in fact be
-$(H_{y}\cap B_{\varepsilon})+y\subset\partial\Omega.$
+Let $\mathcal{F}$ be the subset of $\operatorname{Gr}_{p}(\mathbb{R}^{d})$ appearing in the assumptions of Proposition B.
+For each $n$, there exists $F_{n}\in\mathcal{F}$ such that $g_{n}$ does not contract $F_{n}$, and set $F_{n}:=L(g_{n})\cdot F_{n}$.
+By the non-contracting assumption, we know that $L(g_{n})\cdot B_{\varepsilon}$ contains $F_{n}\cap B_{\varepsilon}$.
+It follows from (1) that $\Omega$ contains $y_{n}+(F_{n}\cap B_{\varepsilon})$, establishing the second point.
+By compactness of $\mathcal{F}$, we may assume (up to extracting a subsequence) that $F_{n}$ converges in the Grassmannian to some $F^{y}\in\mathcal{F}$.
+Then $\overline{\Omega}$ contains $y+(F^{y}\cap B_{\varepsilon})$, and since $y\in\partial\Omega$, the inclusion must in fact be
+$(F^{y}\cap B_{\varepsilon})+y\subset\partial\Omega.$
 This proves the third point and completes the proof. ∎
 Remark 3.3.
-In the previous lemma, we do not use the transversality of the elements of the closure of $\mathcal{F}$.
+In the previous lemma, we do not use the transversality of the elements of $\mathcal{F}$.
 In particular, the lemma asserts that each point of the boundary contains a small neighborhood of an affine subspace.
-The next step is to show the uniqueness of the subspace $H_{y}$. Before doing so, we record the following elementary fact about transversality.
+The next step is to show the uniqueness of the subspace $F^{y}$. Before doing so, we record the following elementary fact about transversality.
 Lemma 3.4.
@@
 Lemma 3.5.
-Let $y\in\partial\Omega$, and let $H_{y}$ be an element of the closure of $\mathcal{F}$ as in Lemma 3.2.
-Then $H_{y}$ is the unique element of $\overline{\mathcal{F}}$ with the property that there exists a neighborhood $\mathcal{V}$ of $y$ in $\Omega$ such that $(y+H_{y})\cap\mathcal{V}\subset\partial\Omega$.
+Let $y\in\partial\Omega$, and let $F^{y}$ be an element of the closure of $\mathcal{F}$ as in Lemma 3.2.
+Then $F^{y}$ is the unique element of $\mathcal{F}$ with the property that there exists a neighborhood $\mathcal{V}$ of $y$ in $\Omega$ such that $(y+F^{y})\cap\mathcal{V}\subset\partial\Omega$.
 Proof.
-Let $(y_{n})_{n\in\mathbb{N}}$, $(H_{n})_{n\in\mathbb{N}}$, and $H_{y}$ be as in Lemma 3.2. By shrinking $\varepsilon$ if necessary, we may assume that $B(y,\varepsilon)\subset\mathcal{V}$.
-We proceed by contradiction and assume the existence of a subspace $H\in\overline{\mathcal{F}}$, distinct from $H_{y}$, that satisfies the hypothesis of the lemma.
-Then $H_{y}$ and $H$ are transverse.
-Since $y_{n}\to y$ and $H_{n}\to H_{y}$, and using Lemma 3.4, we deduce that, for $n$ large enough, the set $(H_{n}+y_{n})\cap B(y_{n},\varepsilon)$ intersects $(H+y)\cap B(y,\varepsilon)$.
-However, by Lemma 3.2, we have $(H_{n}+y_{n})\cap B(y_{n},\varepsilon)\subset\Omega$, while $(H+y)\cap B(y,\varepsilon)\subset(H+y)\cap\mathcal{V}\subset\partial\Omega$, a contradiction. ∎
+Let $(y_{n})_{n\in\mathbb{N}}$, $(F_{n})_{n\in\mathbb{N}}$, and $F^{y}$ be as in Lemma 3.2. By shrinking $\varepsilon$ if necessary, we may assume that $B(y,\varepsilon)\subset\mathcal{V}$.
+We proceed by contradiction and assume the existence of a subspace $F\in\mathcal{F}$, distinct from $F^{y}$, that satisfies the hypothesis of the lemma.
+Then by assumption $(2)$ of Proposition B, one has $F^{y}+F=\mathbb{R}^{d}$.
+Since $y_{n}\to y$ and $F_{n}\to F^{y}$, and using Lemma 3.4, we deduce that, for $n$ large enough, the set $(F_{n}+y_{n})\cap B(y_{n},\varepsilon)$ intersects $(F+y)\cap B(y,\varepsilon)$.
+However, by Lemma 3.2, we have $(F_{n}+y_{n})\cap B(y_{n},\varepsilon)\subset\Omega$, while $(F+y)\cap B(y,\varepsilon)\subset(F+y)\cap\mathcal{V}\subset\partial\Omega$, a contradiction. ∎
 The next lemma goes a step further by showing that $\partial\Omega$ contains an affine subspace.
 Lemma 3.6.
-Let $y\in\partial\Omega$, and let $H_{y}$ be as in Proposition 3.2. Then $H_{y}+y\subset\partial\Omega$.
+Let $y\in\partial\Omega$, and let $F^{y}$ be as in Proposition 3.2. Then $F^{y}+y\subset\partial\Omega$.
 Proof.
-Let $\mathcal{U}$ be the set of points $z\in H_{y}+y$ such that $(H_{y}+y)\cap B(z,\varepsilon)\subset\partial\Omega$.
-We will show that $\mathcal{U}$ is both open and closed (and nonempty) in $H_{y}+y$, so that $\mathcal{U}=H_{y}+y$, which will complete the proof.
+Let $\mathcal{U}$ be the set of points $z\in F^{y}+y$ such that $(F^{y}+y)\cap B(z,\varepsilon)\subset\partial\Omega$.
+We will show that $\mathcal{U}$ is both open and closed (and nonempty) in $F^{y}+y$, so that $\mathcal{U}=F^{y}+y$, which will complete the proof.
 Clearly, the set $\mathcal{U}$ is closed and contains $y$.
 We now show that $\mathcal{U}$ is also open.
-Let $z\in\mathcal{U}$ and $z^{\prime}\in(H_{y}+y)\cap B(z,\varepsilon/2)$.
-Then $(H_{y}+y)\cap B(z^{\prime},\varepsilon/2)\subset\partial\Omega$.
-Thus, by Lemma 3.5, we have $H_{y}=H_{z^{\prime}}$.
-By the definition of $H_{z^{\prime}}$, this implies $H_{y}\cap B(z^{\prime},\varepsilon)=H_{z^{\prime}}\cap B(z^{\prime},\varepsilon)\subset\partial\Omega,$ hence $z^{\prime}\in\mathcal{U}$.
+Let $z\in\mathcal{U}$ and $z^{\prime}\in(F^{y}+y)\cap B(z,\varepsilon/2)$.
+Then $(F^{y}+y)\cap B(z^{\prime},\varepsilon/2)\subset\partial\Omega$.
+Thus, by Lemma 3.5, we have $F^{y}=F^{z^{\prime}}$.
+By the definition of $F^{z^{\prime}}$, this implies $F^{y}\cap B(z^{\prime},\varepsilon)=F^{z^{\prime}}\cap B(z^{\prime},\varepsilon)\subset\partial\Omega,$ hence $z^{\prime}\in\mathcal{U}$.
 Therefore $\mathcal{U}$ is open, which completes the proof.
@@
 Proof of Proposition B.
-First, we claim that two affine subspaces contained in $\partial\Omega$ whose linear parts lie in $\overline{\mathcal{F}}$ must be parallel.
-By Lemma 3.6, there exists at least one affine subspace contained in $\partial\Omega$ whose linear part belongs to $\overline{\mathcal{F}}$.
-Assume, for contradiction, that there exist two affine subspaces $H$ and $H^{\prime}$ contained in $\partial\Omega$ with distinct linear parts in $\overline{\mathcal{F}}$.
+First, we claim that two affine subspaces contained in $\partial\Omega$ whose linear parts lie in $\mathcal{F}$ must be parallel.
+By Lemma 3.6, there exists at least one affine subspace contained in $\partial\Omega$ whose linear part belongs to $\mathcal{F}$.
+Assume, for contradiction, that there exist two affine subspaces $H$ and $H^{\prime}$ contained in $\partial\Omega$ with distinct linear parts in $\mathcal{F}$.
 Then $H$ and $H^{\prime}$ are transverse, so they intersect at some point $y\in\partial\Omega$, which contradicts Lemma 3.5.
@@
 Next, let $H_{0}$ denote the common linear part of the hyperplanes contained in $\partial\Omega$.
-Since $\Gamma$ preserves $\Omega$, the group $L(\Gamma)$ preserves $H_{0}$.
+Since $\mathsf{H}$ preserves $\Omega$, the group $L(\mathsf{H})$ preserves $H_{0}$.
 Let $x\in\Omega$ and suppose that $(H_{0}+x)$ is not contained in $\Omega$.
@@
 ∎
+3.2. General reduction when the linear part is contained in a semisimple Lie subgroup
+The goal of this section is to apply the general reduction result of Proposition B to a case of interest, namely when the linear part of $\mathsf{H}$ is contained in a semisimple Lie subgroup $G$ of ${\operatorname{SL}}_{d}(\mathbb{R})$; see Corollary 3.7 below.
+Indeed, in this case, the action of $G$ on the Grassmannians of $\mathbb{R}^{d}$ is well understood through the Cartan decomposition of $G$.
+This notion is recalled in next subsection .
+Our goal in Section 3.3 will be to apply Corollary 3.7 to the case where $G={\operatorname{SO}}_{0}(2,2)$.
+3.2.1. Cartan projection
+Consider the subgroup ${\operatorname{SO}}(d)$ of ${\operatorname{SL}}_{d}(\mathbb{R})$ and the Cartan subspace of $\operatorname{Mat}(d,\mathbb{R})$ given by
+$\overline{\mathfrak{a}}^{+}:=\{\operatorname{diag}(\lambda_{1},\dots,\lambda_{d})\mid\lambda_{1}\geq\dots\geq\lambda_{d}\ ,\ \ \sum_{k=1}^{d}\lambda_{k}=0\}.$
+Then ${\operatorname{SL}}_{d}(\mathbb{R})={\operatorname{SO}}(d)\exp(\overline{\mathfrak{a}}^{+}){\operatorname{SO}}(d)$.
+This is called the Cartan decomposition of ${\operatorname{SL}}_{d}(\mathbb{R})$.
+Given an element $g\in{\operatorname{SL}}_{d}(\mathbb{R})$, we can thus write $g=k\exp(X)k^{\prime}$ according to this decomposition.
+The elements $k,k^{\prime}\in K$ are not unique, but the element $X\in\overline{\mathfrak{a}}^{+}$ is.
+It is called the Cartan projection of $g$ and denoted by $\mu(g)$.
+We also consider the maps $\mu_{1},\dots,\mu_{d}:{\operatorname{SL}}_{d}(\mathbb{R})\rightarrow\mathbb{R}$, so that
+$\mu(g)=\operatorname{diag}(\mu_{1}(g),\dots,\mu_{d}(g))\quad\forall g\in{\operatorname{SL}}_{d}(\mathbb{R}).$
+3.2.2. A corollary of Proposition B
+We can now prove the following Corollary 3.7 of Proposition B. For this, we take the notation of Subsection 3.2.1.
+Corollary 3.7.
+Let $\Omega$ be a domain in $\mathbb{R}^{d}$, different from $\mathbb{R}^{d}$, and let $\mathsf{H}\leq{\operatorname{SL}}_{d}(\mathbb{R})\ltimes\mathbb{R}^{d}$ be a group acting properly and cocompactly on $\Omega$.
+Let $G$ be a reductive Lie subgroup of ${\operatorname{SL}}_{d}(\mathbb{R})$ containing the linear part $L(\mathsf{H})$ of $\mathsf{H}$.
+Let
+$p_{0}:=\operatorname{card}\{1\leq p\leq d\mid\mu_{i}(G)\subset\mathbb{R}_{+}\}.$
+Assume that there exists $\frac{d}{2}\leq p\leq p_{0}$ the closure $\mathcal{F}$ of the $G$-orbit of $\operatorname{Span}(e_{1},\dots,e_{p})$ in $\operatorname{Gr}_{p}(\mathbb{R}^{d})$ consists of pairwise transverse subspaces, where $(e_{1},\dots,e_{d})$ is the canonical basis of $\mathbb{R}^{d}$.
+Then there exists $F\in\mathcal{F}$ that is $L(\mathsf{H})$-invariant, and $\Omega$ is foliated by affine subspaces parallel to $F$.
+Proof.
+Since $G$ is a reductive subgroup of ${\operatorname{SL}}_{d}(\mathbb{R})$, there always exists a compact subgroup $K$ of ${\operatorname{SO}}(d)$ such that $G=K\exp(\mu(G))K$.
+Note that this decomposition does not always coincide with the Cartan decomposition of $G$.
+For all $g\in G$, we can write $g=waw^{\prime}$ according to the above Cartan decomposition and the assumption that $p\leq p_{0}$, and $g$ does not contract $(w^{\prime})^{-1}\operatorname{Span}(e_{1},\dots,e_{p})$.
+Then all the conditions of Proposition B are satisfied, and the corollary follows.
+∎
+Remark 3.8.
+Corollary 3.7 gives an concrete examples where the conditions of Proposition B are satisfied, such as when $d=8$ and $L(\mathsf{H})$ is contained in ${\operatorname{SO}}(4,\mathbb{C})$, or for any $d$ if $L(\mathsf{H})$ is contained in an irreducible simple rank-one Lie subgroup of ${\operatorname{SL}}_{d}(\mathbb{R})$.
+However, it does not cover all the cases where Proposition B applies: an asymtotic version of it, which is part of work in progress of the three authors, applies if $L(\mathsf{H})$ is an $\{\alpha_{p}\}$-Anosov subgroup of ${\operatorname{SL}}_{d}(\mathbb{R})$ in the sense of [Lab06, GW12], and if its $\{\alpha_{p}\}$-limit set satisfies an asymptotic version of the non-contracting assumption of Proposition B.
+This contracting assumption is for instance satisfied whenever $L(\mathsf{H})$ is contained in ${\sf{Sp}}(d,\mathbb{R})$ for even $d\geq 4$ and $p=d/2$, or in ${\operatorname{SO}}(p,d-p)$ for $p\geq 2$ and $d-p\geq\max(p,3)$.
+Hence the geometric reduction still holds for these groups but does not follow from Corollary 3.7.
 3.3. Reduction in ${\operatorname{SO}}(2,2)$-case
-In this subsection, we apply Proposition B to the geometric setting of primary interest, namely $({\operatorname{SO}}(2,2)\ltimes\mathbb{R}^{2,2},\mathbb{R}^{2,2})$-geometry.
+In this subsection, we apply Corollary 3.7 to the geometric setting of primary interest, namely $({\operatorname{SO}}(2,2)\ltimes\mathbb{R}^{2,2},\mathbb{R}^{2,2})$-geometry.
 Throughout the paper, we denote by $\mathbb{R}^{2,2}$ the affine space $\mathbb{R}^{4}$ endowed with a bilinear form of signature $(2,2)$.
@@
 For this quadratic form, each vector of the canonical basis $(e_{1},e_{2},e_{3},e_{4})$ of $\mathbb{R}^{4}$ is isotropic.
-Let us consider the Cartan subgroup of ${\operatorname{SO}}_{0}(2,2)$ given in the basis $(e_{1},e_{2},e_{3},e_{4})$ by
-$A=\left\{\mathrm{diag}(\lambda,\mu,\lambda^{-1},\mu^{-1})\mid\lambda,\mu>0\right\},$
-and the subset
-$A^{+}=\left\{\mathrm{diag}(\lambda,\mu,\lambda^{-1},\mu^{-1})\mid\lambda,\mu\geq 1\right\}.$
-Considering the maximal compact subgroup $K:=S(\operatorname{O}(2)\times\operatorname{O}(2))={\operatorname{SO}}_{0}(2,2)\cap{\operatorname{SO}}(4)$ of ${\operatorname{SO}}_{0}(2,2)$, we have a Cartan decomposition ${\operatorname{SO}}_{0}(2,2)=KA^{+}K$, in the sense of Section 3.1.
-The set $A^{+}$ does not contract the isotropic plane $\mathrm{Span}(e_{1},e_{2})$ (with respect to the usual euclidean norm of $\mathbb{R}^{4}$).
-Let $\mathcal{F}$ be a closed ${\operatorname{SO}}_{0}(2,2)$-orbit of isotropic planes.
-Thus, we are in the context of Proposition B, and the proof follows.
+In the notation of Section 3.2.1 and Corollary 3.7, we have
+$\mu({\operatorname{SO}}_{0}(2,2))=\left\{\mathrm{diag}(\lambda,\mu,\lambda^{-1},\mu^{-1})\mid\lambda,\mu>0\right\}.$
+Then $p_{0}=2$.
+Besides, the ${\operatorname{SO}}_{0}(2,2)$-orbit of the totally isotropic $2$-plane $\operatorname{Span}(e_{1},e_{2})$ is transverse, by Proposition 3.9.
+Hence we are in the context of Corollary 3.7, and the proof follows.
 ∎
@@
 Let $P_{0}={\operatorname{Span}}(T_{1},T_{2})$ be the center of $N$ and assume that $\Gamma\cap I\cong\mathbb{Z}$. Then
-  • If $\Gamma\cap P_{0}$ is trivial, then after passing to a subgroup of index $2$ of $\Gamma$, we have $p(\Gamma)\leq{\operatorname{SL}}(2,\mathbb{R})$.
+  • If $\Gamma\cap P_{0}$ is trivial, then after passing to a subgroup of index $2$ of $\Gamma$, we have $p(\Gamma)\leq{\operatorname{SL}}_{2}(\mathbb{R})$.
   • If $\Gamma\cap P_{0}$ is nontrivial, then after passing to a subgroup of index $2$ of $\Gamma$, we may assume that $p(\Gamma)$ lies in
@@
 ${\operatorname{det}}(B)\,s_{0}=s_{0},\quad B\bigl(v_{0}-s_{0}Jw_{2}\bigr)=v_{0},\quad\forall\,B\in p(\Gamma).$ (11)
-If $\Gamma\cap P_{0}$ is trivial, then $s_{0}\neq 0$, and hence ${\operatorname{det}}(B)=1$ for all $B\in p(\Gamma)$, so $p(\Gamma)\leq{\operatorname{SL}}(2,\mathbb{R})$.
+If $\Gamma\cap P_{0}$ is trivial, then $s_{0}\neq 0$, and hence ${\operatorname{det}}(B)=1$ for all $B\in p(\Gamma)$, so $p(\Gamma)\leq{\operatorname{SL}}_{2}(\mathbb{R})$.
 If $\Gamma\cap P_{0}$ is nontrivial, then $s_{0}=0$, and equation (11) reduces to $Bv_{0}=v_{0}$, with $v_{0}\in P_{0}$.
@@
 We define $L:{\operatorname{GL}}_{2}^{+}(\mathbb{R})\ltimes N\to{\operatorname{GL}}_{2}^{+}(\mathbb{R})\ltimes U$ so that $L(\Gamma)$ is the linear part of $\Gamma$.
-Since $\Gamma\cap P_{0}$ is trivial, Lemma 5.5 implies that $p(\Gamma)\leq{\operatorname{SL}}_{2}(\mathbb{R})$, and thus $L(\Gamma)\leq{\operatorname{SL}}_{2}(\mathbb{R})\ltimes U$ (see 3.10 for the definition of $U$).
+Since $\Gamma\cap P_{0}$ is trivial, Lemma 5.5 implies that $p(\Gamma)\leq{\operatorname{SL}}_{2}(\mathbb{R})$, and thus $L(\Gamma)\leq{\operatorname{SL}}_{2}(\mathbb{R})\ltimes U$ (see 3.12 for the definition of $U$).
 However, elements of ${\operatorname{SL}}_{2}(\mathbb{R})$ commute with $U$, and therefore $L(\Gamma)$ is contained in the direct product ${\operatorname{SL}}_{2}(\mathbb{R})\times U$.